### v1

# Forecasting Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis

Abstract
The forecasting of multivariate urban data presents a complex challenge due to the intricate dependencies between various urban metrics such as weather, air pollution, carbon intensity, and energy demand. This paper introduces a novel multivariate time-series forecasting model that utilizes advanced Graph Neural Networks (GNNs) to capture spatial dependencies among different time-series variables. The proposed model incorporates a decomposition-based preprocessing step, isolating trend, seasonal, and residual components to enhance the accuracy and interpretability of forecasts. By leveraging the dynamic capabilities of GNNs, the model effectively captures interdependencies and improves the forecasting performance. Extensive experiments on real-world datasets, including electricity usage, weather metrics, carbon intensity, and air pollution data, demonstrate the effectiveness of the proposed approach across various forecasting scenarios. The results highlight the potential of the model to optimize smart infrastructure systems, contributing to energy-efficient urban development and enhanced public well-being.
Keywords Multivariate time-series forecasting, Spatio-temporal graph neural networks, Attention mechanism, Time-series decomposition
1 Introduction
Analyzing patterns and predicting trends in urban data, such as electricity demand Nti et al. (2020), air quality Huang et al. (2021); Du et al. (2024), weather and climate Mouatadid et al. (2024), and carbon intensity Maji et al. (2022), is critical for a wide range of applications, from infrastructure planning and operation to disaster prevention, to addressing social inequalities and improving quality of life. Urban data is normally captured by a number of multimodal sensors installed in a large geographical area. Due to the diverse nature of environments in which the sensors are installed, variations in sensor types and specifications, and complex spatial-temporal dependencies within the data, designing accurate prediction models that can be transferred to various domains is not trivial.
Although multivariate time-series forecasting has been extensively studied in the literature, achieving accurate long-term predictions–specifically for horizons ranging from 2 to 6 weeks–remains a significant challenge compared to short- and medium-term forecasting. This difficulty arises from increased uncertainty, changing dynamics, and accumulation of errors. Furthermore, when data is sourced from multiple sensors with diverse modalities, there is currently no universally effective solution that consistently outperforms other models in different domains.
Physics-based dynamical models are either unavailable or inaccurate for long-term predictions. Traditional statistical methods, recurrent neural networks, and Transformer-based models often fail to capture the complexities of multivariate time-series data, leading to large errors in long-term forecasts. In recent years, foundational models have shown promise in time-series forecasting due to their ability to model complex temporal dependencies and handle multivariate data. However, their effectiveness depends on data availability and they may not generalize well to different time-series data, requiring extensive fine-tuning in some cases. Graph Neural Networks (GNNs) offer a promising solution by effectively modeling temporal and spatial dependencies and improving forecast accuracy Wu et al. (2020). This paper explores the application of GNNs to address the challenges of long-term multivariate time-series forecasting in domains such as energy and transportation. Using these advanced models, we aim to improve the accuracy and reliability of the forecast, where reliability refers to the consistency and robustness of predictions across varying conditions. This, in turn, facilitates the optimization of smart infrastructure systems, which are crucial for developing energy-efficient cities and improving public well-being.
Among GNN models, the Graph Attention Network (GAT) Veličković et al. (2017) introduces an attention mechanism to compute edge weights between neighboring nodes based on node features. This improves neighbor information aggregation and enhances GNN performance in various tasks. However, once trained, GAT’s attention mechanism remains fixed, limiting its ability to adapt to evolving graph structures or changing relationships. Building on this foundation, an improved variant called GATv2 Brody et al. (2021) was developed to address the limitations of the original GAT. GATv2 introduces a more expressive attention mechanism that conditions attention scores on both the query and key nodes, enabling it to better capture complex node interactions. Although this enhancement improves flexibility during training, the learned attention weights of the model remain unchanged after training, similar to GAT. This makes GATv2 particularly effective for tasks requiring richer relational modeling, such as multivariate time-series forecasting.
We propose a novel multivariate time-series forecasting model that takes advantage of GATv2’s power to learn spatial dependencies among the variates and Temporal Convolutional Network (TCN) Bai et al. (2018) to learn temporal dependencies within each variate. Our approach involves constructing a graph where nodes represent different variates and edges represent relationships inferred from historical data in an offline mode. Using GATv2, our model dynamically adjusts the attention given to each node’s neighbors, effectively capturing interdependencies and improving forecast accuracy. We further enhance our model by incorporating a pre-processing step that decomposes the time-series data into its fundamental components: trend, seasonal, and residual. This decomposition allows our model to focus on the inherent characteristics of the data, facilitating more accurate forecasts. Extensive evaluation carried out on different urban datasets demonstrates the superior performance of our approach compared to state-of-the-art time-series forecasting models, highlighting its potential for practical applications in various domains.
The contributions of this paper are summarized below:
  • We propose a novel multivariate time-series forecasting model that uses GATv2 and TCN to learn the salient spatial and temporal features that enhance the accuracy of long-term forecasts.
  • We implement a decomposition-based preprocessing step to isolate the trend, seasonal, and residual components of the time series, improving forecast accuracy.
  • We validate our model’s effectiveness through extensive evaluation on diverse real-world datasets, showcasing significant improvements over the state-of-the-art in time-series forecasting across these datasets.
It is worth noting that, in some applications, forecasts related to multiple types of urban data are often integrated into optimization or decision-making problems to minimize costs, energy consumption, and carbon emissions. For example, forecasts of electricity consumption, weather, and carbon intensity must be incorporated simultaneously to determine how to distribute computing jobs across data centers located in different geographical areas, thereby minimizing costs and carbon emissions Radovanović et al. (2023). Thus, even a slight improvement in the prediction accuracy in each dataset could translate to considerable savings in applications that rely on multiple datasets, highlighting the importance of this contribution.
2 Related Work
Classical statistical methods have been the cornerstone of time-series forecasting for decades. Models such as the Autoregressive Integrated Moving Average (ARIMA) Box and Jenkins (1968); Box and Pierce (1970) and its variants have historically been used due to their simplicity and interpretability. ARIMA models work by modeling the autocorrelation in the data, making them suitable for univariate time-series forecasting. Another well-known method is Exponential Smoothing (ETS) Holt (2004); Winters (1960), which assigns exponentially decreasing weights to past observations. Despite their widespread use, these classical methods have notable limitations. They typically assume stationarity and linearity in the data, which can make them unsuitable for more complex, nonlinear, and large-scale multivariate time-series datasets. Furthermore, as the number of variables increases, these models often struggle with scalability and may overfit data. This underscores the need for more advanced techniques that can handle the intricacies of urban time-series data.
Machine learning has brought new techniques to time-series forecasting. Traditional models like Support Vector Machines (SVM) Boser et al. (1992), Decision Trees, and Random Forests Breiman (2001) have been adapted to predict future values based on historical data. These models are capable of capturing nonlinear relationships better than classical statistical methods. Ensemble methods such as Gradient Boosting Machines (GBM) Friedman (2001) and XGBoost Chen and Guestrin (2016) have been particularly successful in combining the predictions of multiple models to improve accuracy. However, these models often require significant feature engineering and may not fully capture the temporal dependencies inherent in time-series data.
Deep learning has revolutionized time-series forecasting by making it possible to learn nonlinear and complex temporal dependencies directly from the data. Early deep learning models for time-series forecasting, such as Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) networks and Gated Recurrent Units (GRUs) Chung et al. (2014), focused on capturing long-term dependencies through recurrent neural network (RNN) architectures.
More recent advances include models such as N-BEATS Oreshkin et al. (2019), which employs backward and forward residual links and a deep stack of fully connected layers to achieve state-of-the-art performance on various benchmark datasets. Transformer-based models, such as Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and Fedformer Zhou et al. (2022), have also shown significant promise by leveraging self-attention mechanisms to capture long-range dependencies more effectively than RNN-based approaches. However, some recent studies show that transformer-based models may not always be the best fit for time-series forecasting Zeng et al. (2023). Transformer-based models may struggle with effectively capturing the inherent temporal patterns in time-series data due to their reliance on self-attention mechanisms. DLinear and NLinear models Zeng et al. (2023) take a simpler and more efficient approach that mitigates these issues by focusing on linear models that are more adept at handling the sequential nature of time-series data.
One recent related work is CycleNet Lin et al. (2024), which introduces Residual Cycle Forecasting (RCF) to model periodic patterns using learnable recurrent cycles and predict residual components. However, CycleNet assumes stable periodicities, making it less effective for dynamic or irregular cycle datasets. Additionally, it models each time-series channel independently, failing to capture interdependencies between variables, which can limit its performance on multivariate time-series data.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for time-series forecasting, particularly for multivariate time-series data. GNNs are designed to capture spatial dependencies by representing the data as graphs, where nodes represent variables, and edges represent dependencies between these variables. One notable approach is the MTGNN Wu et al. (2020) framework, which integrates temporal and graph convolution modules to capture both temporal and spatial dependencies, using a graph learning layer to adaptively construct an adjacency matrix that reflects hidden relationships among time-series variables. Another example is the Multivariate Time-Series Graph Ordinary Differential Equation (MTGODE) Jin et al. (2022) framework, which uses graph neural networks to model complex interactions among time-series and employs neural ODEs to continuously evolve the data over the graph structure, enhancing forecasting accuracy and robustness.
Defining optimal graph structures and integrating graph learning with time-series forecasting in an end-to-end manner remains a significant challenge. While recent advances have introduced adaptive graph learning Zhang et al. (2020) and automated spatio-temporal fusion techniques Li and Zhu (2021), these methods come with their own limitations. Specifically, adaptive graph learning can be computationally expensive, and spatio-temporal fusion mechanisms may struggle to effectively capture long-range dependencies, leading to potential inaccuracies in forecasting Kumar et al. (2024).
Recent advancements in time-series forecasting feature foundation models such as TimesFM Das et al. (2023) and MOIRAI Woo et al. (2024), which generalize across diverse datasets without additional training. TimesFM, using a decoder-only architecture with input patching, achieves near state-of-the-art zero-shot performance. MOIRAI, with a masked encoder architecture, addresses challenges like cross-frequency learning and any-variate forecasting, leveraging the Large-scale Open Time-Series Archive (LOTSA) Woo et al. (2024) for training. Both models demonstrate the potential of foundation models in handling complex and varied time-series forecasting tasks. However, these models can be computationally intensive, require significant resources for training and deployment, and often lack interpretability.
Overall, while each category of time-series forecasting methods has distinct limitations, the integration of graph neural networks with deep learning techniques represents a promising direction to advance the state-of-the-art in multivariate time-series forecasting.
3 Problem Formulation
Let $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ represent a time-series of $T$ observations, each being $d$-dimensional. Time-series forecasting concerns predicting future values based on past observations. Given a look-back window of length $l$ ending at time $t$, denoted as $X_{t,l}=(x_{t-l+1},\ldots,x_{t-1},x_{t})$, the goal is to forecast the next $h$ steps of the time-series. This is achieved by learning the function $f_{\theta}(X_{t,l})=(\hat{x}_{t+1},\ldots,\hat{x}_{t+h})$, where $\theta$ represents the parameters of the forecasting model. We refer to each pair of look-back and forecast windows as a sample.
4 Methodology
We model the input and output time-series as a graph. Each sensor channel $i$ is represented as a node $v_{i}$ that might have relationships with other channels, represented as directed and weighted edges $e_{ij}$ connecting these nodes. These connections are assumed to be static, i.e., the graph structure $G=(V,E)$ does not change over time. The node features are initially the historical values of the corresponding channels $x_{i}=\{x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{T}\}$, and ultimately the forecast values $\hat{y}_{i}=\{\hat{y}_{i}^{T+1},\hat{y}_{i}^{T+2},\ldots,\hat{y}_{i}^{T+H}\}$.
The proposed methodology has two main parts:
  1. 1.
    Graph Structure Learning: This involves learning the adjacency matrix $A$ capturing the relationships between nodes. The matrix $A\in\mathbb{R}^{D\times D}$ where $D$ is the number of nodes, and each element $a_{ij}$ represents the weight of the edge from node $v_{i}$ to node $v_{j}$.
  2. 2.
    Spatio-Temporal Forecasting: This is the core of our model, responsible for learning and prediction. It comprises three subcomponents:
    1. (a)
      Decomposition: Each input sample is decomposed into its fundamental components to better capture underlying patterns.
    2. (b)
      Spatio-Temporal Feature Extraction: This step extracts features from the graph using both spatial and temporal information. For each node, $v_{i}$, the feature vector $h_{i}$ is derived by combining its historical values $x_{i}$ and the influence of its neighboring nodes as defined by $A$.
    3. (c)
      Forecasting: Finally, the model uses the extracted features to predict future values. The forecast values $\hat{Y}$ for all nodes are given by $\hat{Y}=f(H,A)$, where $H$ is the matrix of all node features and $f$ is the forecasting function.
Each of these parts will be explained in more detail in the subsequent sections.
4.1 Graph structure learning
Since the true underlying graph is unknown, we infer it from the data. To this end, we first decompose the time-series into its fundamental components and then infer the underlying graph for each component independently. This entire process is conducted offline. The graph structure is assumed to be static, hence it is not updated over time.
4.1.1 Decomposition
The first step in our approach involves decomposing the time-series into three fundamental components: trend, residual, and one or more seasonal components. This decomposition is crucial as it isolates the inherent characteristics of the time-series, thus improving the accuracy of forecasting and analysis Sohrabbeig et al. (2023). The decomposition process begins with the application of a moving weighted average filter to the time-series. The weights for this kernel are derived from the tricube function given below:
$\text{tricube}(x)=\begin{cases}(1-|x|^{3})^{3},&\text{if }|x|\leq 1,\\ 0,&\text{if }|x|>1.\end{cases}$ (1)
It assigns higher weights to nearby data points and lower weights to distant ones. The extracted trend is then subtracted from the original time-series to obtain the detrended series.
Next, the period of seasonality (e.g., daily or weekly patterns) must be determined. The period of seasonality can be identified using methods such as autocorrelation analysis, Fourier Transform, data visualization, or expert prior knowledge; in this work, we relied on data visualization and prior expert knowledge. To capture seasonal patterns in the detrended series, an aggregation approach with a sliding window of appropriate length and stride is used. The seasonal pattern is identified by averaging the segments covered by the sliding window. Fig. 1 illustrates this process.
Figure 1: Extracting seasonal patterns using a sliding window of size $l$ and a stride of $m$ times the length of the seasonality period, respectively. In this example, $l{=}3$ and $m{=}1$.
In each step of moving the sliding window, $S_{i}$ is calculated using the following equation:
$S_{i}=\frac{1}{l}\sum_{j=i}^{i+l-1}B_{i},$ (2)
where $B_{i}$ is the i-th block of the detrended time-series. Each block is a vector containing $p$ data points, where $p$ represents the length of the seasonality period, ranging from index $i\times p$ to $(i+1)\times p-1$. To preserve the input length, padding is added to both ends of the time-series. This padding consists of repeated copies of the first and last blocks.
Finally, the so-obtained seasonal components are subtracted from the detrended series to obtain the residual component. Each variate of the multivariate time-series, except for date and time features, undergoes this decomposition process separately. By the end of this procedure, we have extracted a trend, a residual, and one or multiple seasonal components for each variate, ensuring that the unique characteristics of each time-series are accurately captured and ready for the subsequent analysis.
4.1.2 Learning graph structure for each component
After decomposing the time-series into its constructing components, we construct the underlying graph for each component by connecting a subset of nodes. To ensure computational efficiency, each component is downsampled to a level that preserves its general patterns.
Next, the distance between two variates is measured using the Dynamic Time Warping (DTW) Müller (2007) metric. DTW is chosen for its robustness in handling nonlinear alignments, providing reliable similarity measurements between time-series. Finally, we connect each node to its $K$ nearest neighbors according to the DTW distance, with $K$ being a hyperparameter in our approach. The described procedure is outlined in Alg. 1.
Algorithm 1 Time-series Graph Construction
0:  time-series data $\{X_{i}\}_{i=1}^{D}$, seasonality periods $P_{1},P_{2},...,P_{r}$, number of nearest neighbors K
0:  Underlying Graph G
1:  Pattern Extraction:
2:  for each time-series $X_{i}$ do
3:     Extract trend component $T_{i}$ using a moving weighted average with tricube weights and then downsample
4:     Detrend the series: $\hat{X}_{i}=X_{i}-T_{i}$
5:     Extract seasonality component $S_{i}$
6:     for each seasonality period $P$ in $\{P_{1},P_{2},...,P_{r}\}$ do
7:        Apply a moving aggregation approach with window length $l\times P$ and stride $m\times P$
8:        Compute seasonality component $S_{i}$ as the average of the $n$ periods
9:     end for
10:  end for
11:  Distance Metric:
12:  Use DTW metric to calculate the distance between any two pairs of variates and construct distance matrix $D$ as shown in Fig. 3
13:  Construct the Underlying Graph
14:  Connect each node to its K nearest neighbors
Fig. 2 depicts the inferred graph from the trend component of the carbon intensity dataset, described in Section 5.1. Here the inferred graph is directional, as one variate might be dependent on the lagged version of the other one, but the opposite is not necessarily true. Fig. 3 illustrates the Dynamic Time Warping (DTW) distances between different nodes for both the trend and seasonal components of the air pollution dataset. These heatmaps reveal distinct patterns of similarity, with lower normalized DTW values indicating stronger temporal alignment between nodes.
Figure 2: The carbon intensity value is shown for a subset of states in the US at a given point in time. Dark brown (light yellow) indicates higher (lower) carbon intensity in a jurisdiction. Data has been collected from a number of power generation companies, which may supply power to different states. Consequently, the relationship between the states and energy-providing companies is many-to-many. Some arrows extend beyond the map because the image shows only a portion of the full area, and some zones connect to others not included in the display.
Figure 3: Heatmap of DTW distances between different nodes for (a) trend and (b) seasonal components of the air pollution dataset, described in Section 5.1. Note that the distance matrices are symmetrical, with lower values indicating greater closeness. All values are min-max normalized to ensure consistent scaling.
4.2 Spatio-Temporal Forecasting
Figure 4: The overall architecture of proposed forecasting model. After decomposition, each component is handled independently.
Now that we have learned the respective graph structure for each constituent component in an offline fashion, as described in Section 4.1.1, we proceed with the spatio-temporal forecasting step. The overall architecture of the proposed forecasting model is illustrated in Fig. 4. The input of the model is a sample of fixed-length historical values. These historical values include date and time features, along with a number of variates that we would like to predict.
4.2.1 Sample Decomposition
Figure 5: Decomposition of an electricity consumption sample over two weeks. The first subplot displays the original data alongside the sum of the trend and seasonal components, showing their closeness.
First, we need to decompose an input sample, without date and time features, into its constituent components: trend, seasonality, and residual. This process follows the procedure described in Section 4.1.1, but without the need for downsampling. Fig. 5 illustrates the decomposition of a randomly selected sample from the Electricity dataset, described in Section 5.1.
4.2.2 Spatial feature extraction
Our feature extraction model handles each component, that is, trend, seasonal, and residual, independently. For each component, we extract and combine three types of features, namely date-time, spatial, and temporal features.
The date-time features are processed through a linear layer to capture some temporal patterns. Simultaneously, spatial and temporal features are extracted from the input data. These temporal and spatial features are then combined and passed on to the forecasting module that uses them to generate the final output. Fig. 6 is a schematic diagram of the feature extraction module.
$\hat{y}_{t}=f_{\text{forecast}}\left(f_{\text{time}}(x_{\text{time}})\oplus f_{\text{spatio-temporal}}(x)\right)$ (3)
Figure 6: One layer of spatio-temporal feature extraction with time embedding. The input to this module has n channels, $C_{0}$ to $C_{n}-1$. Each component from the decomposition module—trend, seasonal, and residual—is processed independently using this layer. The spatial feature extractor is a Graph Attention Network v2 (GATv2), with a pre-inferred, static underlying graph structure that does not change over time.
Figure 7: Modeling of multivariate time-series as a graph. Each node represents a variate, and its historical data serves as the features of that node
For spatial feature extraction, we utilize a model based on GATv2 Brody et al. (2021). In this model, each variate of the time-series is treated as a node, with its historical values within the sample serving as its features, as depicted in Fig. 7. GATv2 requires a connection graph to understand the relationships between different nodes (variates). The connection graph is inferred using the method described in Section 4.1.
In a GNN, each node iteratively updates its state by aggregating information from its neighbors. The general formulation of a GNN layer is as follows:
$\mathbf{h}_{i}^{\prime}=\sigma\left(\mathbf{W}\mathbf{h}_{i}+\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j}\mathbf{h}_{j}\right)$ (4)
where $\mathbf{h}_{i}$ is the feature vector of node $i$, $\mathcal{N}_{i}$ represents the set of neighbors of node $i$, $\mathbf{W}$ and $\mathbf{W}_{j}$ are learnable weight matrices, and $\sigma$ is a nonlinear activation function. This aggregation allows the model to capture the dependencies between nodes and their neighbors, which is crucial for spatial feature extraction.
In GNNs, Equation 4 is used to calculate the updated node representation as a weighted sum of the neighbors’ representation. To increase learning capacity, we can incorporate this method with an attention mechanism. The Graph Attention Network (GAT) Veličković et al. (2017) enhances the GNN framework by incorporating an attention mechanism. This mechanism allows nodes to assign different amounts of importance to the neighbors, which enables capturing of changing dynamics. Specifically, the attention score $e_{ij}$ between node $i$ and its neighbor $j$ is computed as:
$e_{ij}=\text{LeakyReLU}\left(\mathbf{a}^{T}[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}]\right)$ (5)
where $\mathbf{a}$ is a learnable weight vector, $\mathbf{W}$ is a weight matrix, $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the feature vectors of nodes $i$ and $j$, respectively, and $\parallel$ denotes concatenation. The attention scores are then normalized using Softmax:
$\alpha_{ij}=\text{softmax}_{j}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$ (6)
The final output representation of node $i$ is computed as a weighted sum of its neighbors’ features:
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (7)
This attention mechanism allows GAT to focus on the most relevant neighbors, improving its ability to capture complex relationships in the graph.
GATv2 Brody et al. (2021), builds on the original GAT by addressing its limitations and introducing a more expressive dynamic attention mechanism. In GATv2, the attention score computation is modified to condition both the query and key nodes, enhancing the model’s flexibility. The new attention score $e_{ij}$ is defined as:
$e_{ij}=\mathbf{a}^{T}\text{LeakyReLU}\left(\mathbf{W}[\mathbf{h}_{i}\parallel\mathbf{h}_{j}]\right)$ (8)
In this architecture, we employed just one layer of GATv2. The decision of whether to use one layer of GATv2 or stack more of them has been made empirically. Fig. 9 in Appendix A justifies why one layer of the GATv2 layer was employed for the carbon intensity dataset.
4.2.3 Temporal Feature Extraction
For temporal feature extraction, we utilize Temporal Convolutional Networks (TCNs) Bai et al. (2018). TCNs are designed to capture temporal dependencies in sequential data by leveraging convolution across time steps. The key components of TCNs are causal convolutions, dilations, and residual connections.
Causal convolutions ensure that the model only uses past information to predict future values, preserving the temporal order of the data. The causal convolution for a given layer $l$ is defined as:
$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i}$ (9)
where $\mathbf{y}_{t}^{l}$ is the output at time $t$ for layer $l$, $\mathbf{W}_{i}^{l}$ are the convolutional filters, $\mathbf{x}_{t-i}$ are the inputs from the previous layer, and $k$ is the filter size.
Dilated convolutions expand the receptive field of the network without increasing the number of parameters, allowing the model to capture long-range dependencies. The dilation factor $d$ determines the spacing between the kernel elements. A dilated convolution is defined as:
$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i\cdot d}$ (10)
where $d$ is the dilation factor, $k$ is the filter size, and the other symbols retain their previous meanings.
Residual connections help mitigate the vanishing gradient problem and facilitate the training of deeper networks. The output of each TCN layer is combined with its input through a residual connection:
$\mathbf{y}_{t}=\sigma\left(\mathbf{y}_{t}^{l}+\mathbf{x}_{t}\right)$ (11)
4.2.4 Forecasting
Once features are extracted for each component of the time-series, we employ a linear forecasting model to produce forecasts for the entire forecast horizon at once, utilizing a direct multistep forecasting approach. This choice is motivated by the demonstrated effectiveness of similar methods in previous studies Sohrabbeig et al. (2023); Zeng et al. (2023). Forecasts are generated simultaneously and independently for each component—trend, seasonal, and residual—mirroring the approach used in the feature extraction phase. Within each forecasting component, each variate is forecasted in a univariate fashion. This means that only the spatial and temporal features extracted for each specific variate in the previous step are used for forecasting that variate. Nevertheless, these features are influenced by the features of some other variates due to the way that they are generated.
The input to the forecasting module is the combination of spatio-temporal features and time embeddings. In this work, these features and embeddings are added together before being fed into the forecasting task. The decision to add rather than concatenate the time embeddings with the other extracted features was made based on our experimental results. Fig. 6 shows the input for the forecasting module.
The linear forecasting model operates on input data with the following characteristics:
  • Input Sequence Length ($L$): The length of the input time-series sequence.
  • Prediction Length ($H$): The number of future time steps to be predicted.
  • Number of Channels ($D$): The number of separate time-series channels in the input data.
The input to the model is a tensor $\mathbf{X}$ of shape $(B,D,L)$, where $B$ is the batch size, $D$ is the number of channels, and $L$ is the length of the input sequence. During the forward pass, the model computes the prediction $\mathbf{Y}$ using the input tensor $\mathbf{X}$ and the learnable parameters. This is done through a linear transformation followed by the addition of the bias term:
$\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}+\mathbf{b}$ (12)
where $\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}$ represents the Einstein summation operation defined as:
$\mathbf{Y}_{bdh}=\sum_{l=1}^{L}\mathbf{W}_{dhl}\cdot\mathbf{X}_{bdl}$ (13)
for each batch $b$, channel $d$, and prediction time step $h$. This process allows the model to learn a linear mapping from the input sequence to the predicted sequence, leveraging the entire input sequence to inform each prediction step. This implementation is more efficient than the method used by DLinear and NLinear Zeng et al. (2023) models, which iteratively calculates forecasts for each variate.
5 Result
In this section, we first introduce the datasets and baseline models used for evaluation, then evaluate and compare the performance of our proposed model with the baselines in three different settings.
5.1 Datasets
For comprehensive coverage of critical factors influencing time series data, we select four urban datasets including electricity usage, weather conditions, carbon intensity, and air pollution. These datasets provide a robust benchmark for understanding the pros and cons of different time series forecasting models. More details on these datasets are provided below:
Table 1 summarizes all the main characteristics of the above datasets. With the exception of the air pollution dataset, the other datasets did not require any imputation of missing values. For the air pollution dataset, missing values were filled using the average of records corresponding to the same month, day of the week, and hour. As a preprocessing step, all columns in all datasets were standardized.
Table 1: Statistical details of the real-world datasets used in this study.
Dataset No. Features Length Time Resolution
Electricity 321 26,304 1h
Air Pollution 7 34,970 1h
Carbon 54 26,280 1h
Weather 21 52,696 10min
5.2 Baselines
To evaluate our model, we use three highly effective Transformer models, namely Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022), as our baselines. Additionally, we consider two recent state-of-the-art models, NLinear and DLinear Zeng et al. (2023). We also employed a straightforward baseline: Repeat Last, which replicates the last observation. Our next baseline is TimesFM Das et al. (2023), a foundation model for time series forecasting that leverages a decoder-only attention mechanism with input patching. TimesFM is pre-trained on a large corpus of both real-world and synthetic time series data, enabling it to achieve near state-of-the-art zero-shot performance across various datasets without the need for fine-tuning. In these experiments, TimesFM was used solely for inference and was not fine-tuned.
In this study, we did not compare our results with classical models such as ARIMA and SARIMA, due to the extensive time required for their parameter selection, which significantly slows down the training process. Instead, we based our conclusions on experimental results reported in other studies Wu et al. (2021); Zhou et al. (2021), which have shown that these classical models perform poorly on the long sequence time series forecasting (LSTF) task. Therefore, we opted not to include them in our direct comparisons.
5.3 Experiments
We now evaluate the performance of our model (labelled DST in the tables) and compare it with strong baselines in three different settings. In all experiments, the length of the look-back window is 336, and the learning rate is set to 0.0001. The length of the forecast horizon is denoted in the second column of tables 2, 3 and 4. For our model, we decided to extract only one seasonal component due to the short length of the input. For instance, when the length of the look-back window is 336 hours (equivalent to two weeks), it is logical to extract daily seasonal patterns since multiple daily cycles are present in this period while extracting weekly patterns would be less meaningful considering the limited number of weekly cycles.
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.402 0.322 0.421 0.324 0.441 0.334 0.481 0.382 0.532 0.429 0.745 0.589 4.51 0.62 192 0.422 0.333 0.439 0.340 0.447 0.344 0.483 0.376 0.535 0.431 0.813 0.615 3.87 2.06 336 0.427 0.341 0.447 0.350 0.467 0.358 0.488 0.382 0.542 0.431 0.804 0.623 4.47 2.57 720 0.423 0.347 0.469 0.375 0.472 0.379 0.452 0.342 0.549 0.432 0.827 0.644 9.80 7.46 Carbon 96 0.468 0.501 0.476 0.503 0.467 0.492 0.512 0.539 0.519 0.547 1.128 0.780 -0.21 -1.83 192 0.534 0.545 0.541 0.544 0.540 0.547 0.674 0.621 0.584 0.580 1.243 0.829 1.11 0.37 336 0.589 0.581 0.601 0.595 0.597 0.594 0.703 0.659 0.631 0.629 1.328 0.865 1.34 2.19 720 0.721 0.656 0.731 0.669 0.729 0.661 0.878 0.738 0.723 0.650 1.489 0.927 1.10 0.76 Electricity 96 0.131 0.228 0.149 0.252 0.136 0.232 0.190 0.301 0.189 0.298 1.528 0.927 3.68 1.72 192 0.141 0.241 0.160 0.264 0.167 0.269 0.202 0.313 0.204 0.311 1.521 0.927 11.87 8.71 336 0.153 0.260 0.170 0.279 0.169 0.275 0.210 0.318 0.231 0.336 1.530 0.934 9.47 5.45 720 0.171 0.282 0.204 0.314 0.200 0.298 0.239 0.350 0.246 0.361 1.535 0.942 14.5 5.37 Weather 96 0.178 0.261 0.169 0.251 0.175 0.259 0.364 0.406 0.319 0.329 0.258 0.254 -5.33 -3.98 192 0.234 0.315 0.229 0.300 0.225 0.299 0.436 0.439 0.322 0.354 0.339 0.294 -4.00 -7.14 336 0.381 0.443 0.321 0.403 0.352 0.412 0.470 0.469 0.401 0.391 0.421 0.343 -18.69 -29.15 720 0.549 0.536 0.414 0.465 0.421 0.481 0.467 0.447 0.433 0.444 0.465 0.394 -32.60 -36.04
Table 2: Experimental results on real data in multivariate input, multivariate output setting. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.864 0.703 0.874 0.699 0.872 0.701 0.947 0.714 0.914 0.695 1.579 0.869 0.92 -0.57 192 0.903 0.719 0.926 0.728 0.931 0.732 1.008 0.744 0.973 0.739 1.725 0.926 2.48 1.24 336 0.916 0.722 0.954 0.738 0.955 0.745 1.057 0.792 0.969 0.744 1.799 0.948 3.98 2.17 720 0.945 0.735 0.986 0.746 1.015 0.783 1.063 0.814 1.009 0.751 1.877 0.977 4.16 1.47 Carbon 96 0.232 0.370 0.235 0.375 0.244 0.381 0.731 0.858 0.418 0.561 1.771 1.039 1.28 1.33 192 0.248 0.389 0.273 0.406 0.284 0.412 0.779 0.871 0.445 0.594 1.853 1.077 9.16 4.19 336 0.266 0.410 0.307 0.430 0.305 0.427 0.801 0.896 0.457 0.612 1.919 1.106 12.79 3.98 720 0.282 0.425 0.368 0.476 0.364 0.469 0.818 0.914 0.551 0.687 2.025 1.141 22.53 9.38 Electricity 96 0.148 0.273 0.166 0.292 0.169 0.293 0.194 0.309 0.200 0.291 1.221 0.858 10.84 6.19 192 0.176 0.298 0.189 0.312 0.191 0.320 0.211 0.319 0.209 0.319 1.223 0.857 6.87 4.49 336 0.188 0.310 0.209 0.329 0.204 0.321 0.218 0.330 0.211 0.334 1.297 0.881 7.84 3.43 720 0.214 0.335 0.252 0.371 0.255 0.374 0.240 0.351 0.258 0.359 1.369 0.908 10.83 4.56 Weather 96 0.096 0.206 0.127 0.234 0.131 0.237 0.109 0.212 0.112 0.243 0.130 0.262 11.93 2.83 192 0.378 0.416 0.412 0.425 0.423 0.427 0.401 0.419 0.439 0.461 0.451 0.478 5.74 0.72 336 0.780 0.604 0.834 0.613 0.845 0.621 0.806 0.605 0.812 0.604 0.844 0.620 3.23 0.00 720 1.237 0.797 1.408 0.825 1.392 0.818 1.398 0.816 1.418 0.876 1.470 0.934 11.14 2.33
Table 3: Experimental results on real data in multivariate input, univariate output setting. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.867 0.695 0.874 0.699 0.872 0.701 0.961 0.732 0.917 0.699 1.579 0.869 0.57 0.57 192 0.921 0.724 0.926 0.728 0.931 0.732 1.023 0.760 0.971 0.741 1.725 0.926 0.54 0.55 336 0.950 0.736 0.954 0.738 0.955 0.745 1.084 0.803 1.001 0.747 1.799 0.948 0.42 0.27 720 0.984 0.747 0.986 0.746 1.015 0.783 1.092 0.833 1.013 0.755 1.877 0.977 0.20 -0.13 Carbon 96 0.216 0.357 0.235 0.375 0.244 0.381 0.615 0.796 0.422 0.567 1.771 1.039 8.08 4.80 192 0.264 0.398 0.273 0.406 0.284 0.412 0.631 0.818 0.446 0.601 1.853 1.077 3.29 1.97 336 0.293 0.422 0.307 0.430 0.305 0.427 0.722 0.846 0.463 0.615 1.919 1.106 3.3 1.17 720 0.360 0.471 0.368 0.476 0.364 0.469 0.779 0.875 0.557 0.690 2.025 1.141 1.10 -0.42 Electricity 96 0.154 0.277 0.166 0.292 0.169 0.293 0.198 0.311 0.209 0.291 1.221 0.858 7.23 4.81 192 0.173 0.297 0.189 0.312 0.191 0.320 0.214 0.323 0.223 0.342 1.223 0.857 8.46 4.81 336 0.196 0.315 0.209 0.329 0.204 0.321 0.231 0.348 0.231 0.345 1.297 0.881 3.92 1.87 720 0.238 0.357 0.252 0.371 0.255 0.374 0.288 0.389 0.307 0.398 1.369 0.908 5.55 3.77 Weather 96 0.119 0.228 0.127 0.234 0.131 0.237 0.138 0.245 0.157 0.282 0.130 0.262 6.30 2.56 192 0.421 0.433 0.412 0.425 0.423 0.427 0.476 0.501 0.470 0.483 0.451 0.478 -2.18 -1.88 336 0.843 0.621 0.834 0.613 0.845 0.621 0.864 0.658 0.872 0.684 0.844 0.620 -1.07 -1.30 720 1.414 0.826 1.408 0.825 1.392 0.818 1.412 0.835 1.555 0.954 1.470 0.934 -1.58 -0.97
Table 4: Experimental results on real data in univariate input, univariate output. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
As shown in Tables 2, 3, and 4, our proposed model has achieved the best or near-best results across all settings. We present detailed results in three different settings to ensure a fair comparison of the strengths and weaknesses of the baseline models and to clearly explain the effectiveness of our proposed model.
Figure 8: MSE comparison of our proposed model against baselines on different datasets. This is a multivariate input, multivariate output task, with input and output lengths of 336 and 96, respectively. Note that the axis scales are exaggerated for better visualization of the differences between models.
DLinear and NLinear achieve the best performance among the baselines. Despite their impressive performance, they are univariate models that lack a mechanism to extract spatial dependencies. This limitation becomes evident in the multivariate input and univariate forecasting setting, where forecasting a variate may benefit from capturing lagged dependencies with other variates. Univariate models like DLinear or NLinear cannot leverage this information, leading to inferior performance compared to models that can, such as our proposed model and transformer-based models. This is particularly evident in the forecasting of PM2.5 with multivariate input.
Transformer-based forecasting models, such as Autoformer, Fedformer, and Informer, are highly complex and over-parameterized, making them susceptible to overfitting. This issue is apparent in the forecasting of the carbon dataset. Additionally, we have included the inference results of TimesFM, which were obtained without additional training or fine-tuning. This lack of fine-tuning likely contributes to its suboptimal performance, as it might not have adapted well to the specific characteristics of the datasets. Pretrained models are generally not tailored to the nuances of new datasets unless they undergo further training to update their parameters. Consequently, the model’s inability to learn and adjust to the unique temporal patterns and data distribution of the datasets has likely resulted in less accurate forecasts compared to the other models that were specifically trained.
Figure 8 presents a comparative analysis of the Mean Squared Error (MSE) performance of our proposed model, DST, against several baseline models across four datasets: Beijing PM2.5, Carbon, Electricity, and Weather. This evaluation is conducted on a multivariate input-multivariate output forecasting task, where the input and output lengths are set to 336 and 96, respectively. The results demonstrate that DST consistently achieves lower MSE values compared to the baselines, indicating its superior forecasting capability. Notably, traditional transformer-based models, such as Autoformer and Informer, tend to exhibit higher variance and median errors, suggesting increased instability in their predictions. DLinear, a linear baseline, performs competitively but is outperformed by DST, particularly on the Carbon dataset. The y-axis scales are adjusted for better visualization of performance differences among the models
6 Ablations
To thoroughly examine the properties of our proposed model, we conducted ablation studies to evaluate the effectiveness of each of its key modules. The modules under evaluation include (1) Decomposition, (2) Date and Time Embedding, (3) Temporal Feature Extraction, and (4) Spatial Feature Extraction. For each dataset, we compare the model’s performance with and without each of these modules.
Due to the modular and pipelined structure of the proposed architecture, we can easily remove individual modules without impacting the functionality of the remaining components. The only exception is the Date and Time Embedding module, whose output is concatenated with the extracted spatio-temporal features. To simulate the removal of this module, we replace its output with zero-vectors, thereby mitigating its effect on the model’s performance.
Table 5 demonstrates the effectiveness of each module within the model architecture. Among all the components, the temporal feature extraction module plays the most influential role, as indicated by the highest increase in Mean Squared Error (MSE) when this module is removed. This suggests that capturing temporal dependencies is crucial for accurately modeling the datasets, particularly for data characterized by strong temporal patterns, such as air pollution, carbon emissions, and electricity usage. The ablation results also reveal that other modules, such as decomposition, date and time embedding, and spatial feature engineering, contribute significantly to model performance, albeit to a lesser extent. These components work synergistically to enhance predictive accuracy by providing complementary information that captures different aspects of the data, such as spatial correlations and temporal variations. Overall, the study reveals the necessity of taking a multi-faceted approach that integrates various techniques to effectively handle complex, real-world datasets.
Additionally, we conducted experiments to evaluate the impact of the number of GATv2 layers. Across all experiments, a single layer for spatial feature extraction consistently yielded the best performance. Figure 9 illustrates how the MSE changes as more GATv2 layers are added for the carbon intensity dataset.
Dataset Air Pollution Carbon Electricity Weather
w/o Decomposition 0.441 0.474 0.154 0.179
w/o Date&Time Emb. 0.424 0.482 0.149 0.158
w/o Temporal F.E. 0.452 0.491 0.162 0.161
w/o Spatial F.E. 0.439 0.488 0.156 0.165
Original 0.404 0.460 0.133 0.155
Table 5: Ablation studies on the effectiveness of key modules. Reported values are the MSE of each setting.
7 Conclusion
In this paper, we introduced a novel multivariate time-series forecasting model that leverages the advanced capabilities of GATv2 to capture spatial dependencies among different time-series variables. Our approach, which incorporates a decomposition-based preprocessing step, has demonstrated incremental improvements in forecast accuracy and reliability across various real-world datasets. The experimental results underscore the potential of our model to outperform existing state-of-the-art methods, particularly in complex domains such as energy management, weather prediction, carbon intensity forecasting, and air pollution monitoring. By effectively modeling the intricate interdependencies within multivariate time-series data, our proposed model provides a robust framework for long-term forecasting, contributing to the development of smart infrastructure systems and enhancing societal outcomes. Future work will focus on further optimizing the model, exploring its applicability to additional domains, and integrating adaptive graph learning techniques to enhance its performance in dynamic environments.
Appendix
Appendix A Number of GATv2 layers
Figure 9: Changes of MSE loss by increasing the number of GATv2 layers used for spatial feature extraction on carbon intensity dataset.

### v2

@@
-Forecasting Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis
+Long-Term Forecasting of Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis
 Abstract
-The forecasting of multivariate urban data presents a complex challenge due to the intricate dependencies between various urban metrics such as weather, air pollution, carbon intensity, and energy demand.
-This paper introduces a novel multivariate time-series forecasting model that utilizes advanced Graph Neural Networks (GNNs) to capture spatial dependencies among different time-series variables.
-The proposed model incorporates a decomposition-based preprocessing step, isolating trend, seasonal, and residual components to enhance the accuracy and interpretability of forecasts.
-By leveraging the dynamic capabilities of GNNs, the model effectively captures interdependencies and improves the forecasting performance.
-Extensive experiments on real-world datasets, including electricity usage, weather metrics, carbon intensity, and air pollution data, demonstrate the effectiveness of the proposed approach across various forecasting scenarios.
-The results highlight the potential of the model to optimize smart infrastructure systems, contributing to energy-efficient urban development and enhanced public well-being.
+Long-term forecasting of multivariate urban data poses a significant challenge due to the complex spatiotemporal dependencies inherent in such datasets.
+This paper presents DST, a novel multivariate time-series forecasting model that integrates graph attention and temporal convolution within a Graph Neural Network (GNN) to effectively capture spatial and temporal dependencies, respectively.
+To enhance model performance, we apply a decomposition-based preprocessing step that isolates trend, seasonal, and residual components of the time series, enabling the learning of distinct graph structures for different time-series components.
+Extensive experiments on real-world urban datasets—including electricity demand, weather metrics, carbon intensity, and air pollution—demonstrate the effectiveness of DST across a range of forecast horizons, from several days to one month.
+Specifically, our approach achieves an average improvement of 2.89% to 9.10% in long-term forecasting accuracy over state-of-the-art time-series forecasting models.
 Keywords Multivariate time-series forecasting, Spatio-temporal graph neural networks, Attention mechanism, Time-series decomposition
@@
 Analyzing patterns and predicting trends in urban data, such as electricity demand Nti et al.
-(2020), air quality Huang et al.
+(2020), traffic flow and air quality Huang et al.
 (2021); Du et al.
-(2024), weather and climate Mouatadid et al.
+(2024), weather and climate variables Mouatadid et al.
 (2024), and carbon intensity Maji et al.
-(2022), is critical for a wide range of applications, from infrastructure planning and operation to disaster prevention, to addressing social inequalities and improving quality of life.
-Urban data is normally captured by a number of multimodal sensors installed in a large geographical area.
-Due to the diverse nature of environments in which the sensors are installed, variations in sensor types and specifications, and complex spatial-temporal dependencies within the data, designing accurate prediction models that can be transferred to various domains is not trivial.
-Although multivariate time-series forecasting has been extensively studied in the literature, achieving accurate long-term predictions–specifically for horizons ranging from 2 to 6 weeks–remains a significant challenge compared to short- and medium-term forecasting.
-This difficulty arises from increased uncertainty, changing dynamics, and accumulation of errors.
-Furthermore, when data is sourced from multiple sensors with diverse modalities, there is currently no universally effective solution that consistently outperforms other models in different domains.
-Physics-based dynamical models are either unavailable or inaccurate for long-term predictions.
-Traditional statistical methods, recurrent neural networks, and Transformer-based models often fail to capture the complexities of multivariate time-series data, leading to large errors in long-term forecasts.
+(2022), is critical for a wide range of applications, including infrastructure planning, disaster preparedness, and addressing social inequalities.
+Urban data is typically multimodal and collected across large geographical areas.
+Due to variations in sensor types and deployment environments, and complex spatiotemporal dependencies within the data, building accurate and transferrable forecasting models is a non-trivial task.
+While multivariate time-series forecasting has been widely studied, achieving accurate long-term predictions, specifically over horizons of two to six weeks, remains a significant challenge Zhou et al.
+(2021).
+This difficulty arises from the accumulation of errors and shifting dynamics.
+Furthermore, when data originates from multiple heterogeneous sensors, there is currently no universally effective model that consistently performs well across domains.
+Physics-based dynamical models, where available, are inaccurate for long-term forecasting Mouatadid et al.
+(2024).
+Traditional statistical methods, recurrent neural networks, and transformer-based models often fail to capture the complex dependencies inherent in multivariate time-series data, leading to performance degredation on long-term forecasts.
 In recent years, foundational models have shown promise in time-series forecasting due to their ability to model complex temporal dependencies and handle multivariate data.
-However, their effectiveness depends on data availability and they may not generalize well to different time-series data, requiring extensive fine-tuning in some cases.
-Graph Neural Networks (GNNs) offer a promising solution by effectively modeling temporal and spatial dependencies and improving forecast accuracy Wu et al.
+However, their success hinges on access to large-scale pretraining data Woo et al.
+(2024), and they often require extensive fine-tuning to adapt to new domains, limiting their practical applicability.
+Graph Neural Networks (GNNs) offer a promising alternative by naturally capturing both temporal and spatial dependencies, thereby improving forecast accuracy in multivariate settings Wu et al.
 (2020).
-This paper explores the application of GNNs to address the challenges of long-term multivariate time-series forecasting in domains such as energy and transportation.
-Using these advanced models, we aim to improve the accuracy and reliability of the forecast, where reliability refers to the consistency and robustness of predictions across varying conditions.
-This, in turn, facilitates the optimization of smart infrastructure systems, which are crucial for developing energy-efficient cities and improving public well-being.
-Among GNN models, the Graph Attention Network (GAT) Veličković et al.
-(2017) introduces an attention mechanism to compute edge weights between neighboring nodes based on node features.
-This improves neighbor information aggregation and enhances GNN performance in various tasks.
-However, once trained, GAT’s attention mechanism remains fixed, limiting its ability to adapt to evolving graph structures or changing relationships.
-Building on this foundation, an improved variant called GATv2 Brody et al.
-(2021) was developed to address the limitations of the original GAT.
-GATv2 introduces a more expressive attention mechanism that conditions attention scores on both the query and key nodes, enabling it to better capture complex node interactions.
-Although this enhancement improves flexibility during training, the learned attention weights of the model remain unchanged after training, similar to GAT.
-This makes GATv2 particularly effective for tasks requiring richer relational modeling, such as multivariate time-series forecasting.
-We propose a novel multivariate time-series forecasting model that takes advantage of GATv2’s power to learn spatial dependencies among the variates and Temporal Convolutional Network (TCN) Bai et al.
-(2018) to learn temporal dependencies within each variate.
-Our approach involves constructing a graph where nodes represent different variates and edges represent relationships inferred from historical data in an offline mode.
-Using GATv2, our model dynamically adjusts the attention given to each node’s neighbors, effectively capturing interdependencies and improving forecast accuracy.
-We further enhance our model by incorporating a pre-processing step that decomposes the time-series data into its fundamental components: trend, seasonal, and residual.
-This decomposition allows our model to focus on the inherent characteristics of the data, facilitating more accurate forecasts.
-Extensive evaluation carried out on different urban datasets demonstrates the superior performance of our approach compared to state-of-the-art time-series forecasting models, highlighting its potential for practical applications in various domains.
+However, the underlying graph structure is typically unknown and must be inferred from the data.
+In this paper, we explore the use of GNNs to address the challenges of long-term forecasting of multivariate urban time-series data, with an emphasis on improving forecast accuracy.
+A key insight of our approach is that the latent graph structures corresponding to different time-series components (e.g., trend, seasonality) are often distinct.
+This observation motivates the application of time-series decomposition prior to graph structure learning.
+We propose a novel model that combines the Graph Attention Network (GAT) Veličković et al.
+(2017) for learning spatial dependencies among variates with a Temporal Convolutional Network (TCN) Bai et al.
+(2018) to capture temporal dependencies within each variate.
+We construct a graph where nodes represent individual variates and edges encode dependencies inferred from historical data.
+Leveraging GATv2 Brody et al.
+(2021), the model dynamically adjusts the influence of each node on neighboring nodes.
+To further enhance accuracy, we introduce a preprocessing step that decomposes the time-series into its fundamental components: trend, seasonal, and residual components.
+This enables the model to focus on the intrinsic structure of the data, using a distinct graph structure for each component.
+We evaluate our model on multiple real-world urban datasets, demonstrating significant improvements in long-term forecasting performance over the state-of-the-art.
+These results underscore the model’s potential for practical deployment in diverse application domains.
 The contributions of this paper are summarized below:
-  • We propose a novel multivariate time-series forecasting model that uses GATv2 and TCN to learn the salient spatial and temporal features that enhance the accuracy of long-term forecasts.
-  • We implement a decomposition-based preprocessing step to isolate the trend, seasonal, and residual components of the time series, improving forecast accuracy.
-  • We validate our model’s effectiveness through extensive evaluation on diverse real-world datasets, showcasing significant improvements over the state-of-the-art in time-series forecasting across these datasets.
-It is worth noting that, in some applications, forecasts related to multiple types of urban data are often integrated into optimization or decision-making problems to minimize costs, energy consumption, and carbon emissions.
-For example, forecasts of electricity consumption, weather, and carbon intensity must be incorporated simultaneously to determine how to distribute computing jobs across data centers located in different geographical areas, thereby minimizing costs and carbon emissions Radovanović et al.
+  • We propose a novel multivariate time-series forecasting model, called DST, that integrates GATv2 and TCN to learn the salient spatial and temporal features, thereby enhancing the accuracy of long-term forecasting.
+  • We implement a decomposition-based preprocessing step to isolate trend, seasonal, and residual components of the time-series, and learn a graph structure for each component independently.
+Together, they allow the model to tailor its predictions to distinct temporal dynamics, significantly improving its accuracy.
+  • We validate our model’s effectiveness through extensive evaluation on four real-world datasets, showcasing on average 2.89% to 9.10% improvement over the state-of-the-art in time-series forecasting across these datasets.
+We note that accurate forecasts across different types of urban data are often integral to decision-making tasks aimed at minimizing cost, energy consumption, and carbon emissions.
+For example, effectively distributing workloads across geographically dispersed datacenters requires simultaneous predictions of electricity demand, weather, and carbon intensity Radovanović et al.
 (2023).
-Thus, even a slight improvement in the prediction accuracy in each dataset could translate to considerable savings in applications that rely on multiple datasets, highlighting the importance of this contribution.
+Thus, even a slight improvement in the prediction accuracy in each dataset could translate to considerable savings in applications that rely on multiple such datasets.
+This highlights the broader impact of our contributions.
 2 Related Work
-Classical statistical methods have been the cornerstone of time-series forecasting for decades.
-Models such as the Autoregressive Integrated Moving Average (ARIMA) Box and Jenkins (1968); Box and Pierce (1970) and its variants have historically been used due to their simplicity and interpretability.
-ARIMA models work by modeling the autocorrelation in the data, making them suitable for univariate time-series forecasting.
-Another well-known method is Exponential Smoothing (ETS) Holt (2004); Winters (1960), which assigns exponentially decreasing weights to past observations.
+Statistical methods have been the cornerstone of time-series forecasting for decades.
+Time-series models such as the Autoregressive Integrated Moving Average (ARIMA) Box and Pierce (1970) and its variants have historically been used due to their simplicity and interpretability.
+Another widely used method is Exponential Smoothing (ETS) Holt (2004); Winters (1960), which assigns exponentially decreasing weights to past observations.
 Despite their widespread use, these classical methods have notable limitations.
@@
 Furthermore, as the number of variables increases, these models often struggle with scalability and may overfit data.
-This underscores the need for more advanced techniques that can handle the intricacies of urban time-series data.
-Machine learning has brought new techniques to time-series forecasting.
-Traditional models like Support Vector Machines (SVM) Boser et al.
-(1992), Decision Trees, and Random Forests Breiman (2001) have been adapted to predict future values based on historical data.
-These models are capable of capturing nonlinear relationships better than classical statistical methods.
+Traditional machine learning models such as Support Vector Machines (SVM), Decision Trees, and Random Forests have been adapted to predict future values based on historical data.
+These models are capable of capturing nonlinear relationships better than statistical methods.
 Ensemble methods such as Gradient Boosting Machines (GBM) Friedman (2001) and XGBoost Chen and Guestrin (2016) have been particularly successful in combining the predictions of multiple models to improve accuracy.
@@
 Deep learning has revolutionized time-series forecasting by making it possible to learn nonlinear and complex temporal dependencies directly from the data.
-Early deep learning models for time-series forecasting, such as Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) networks and Gated Recurrent Units (GRUs) Chung et al.
-(2014), focused on capturing long-term dependencies through recurrent neural network (RNN) architectures.
+Early deep learning models for time-series forecasting
+focused on capturing long-term dependencies through recurrent neural network (RNN) architectures.
 More recent advances include models such as N-BEATS Oreshkin et al.
@@
 (2023).
-Transformer-based models may struggle with effectively capturing the inherent temporal patterns in time-series data due to their reliance on self-attention mechanisms.
+Particularly, they may struggle with effectively capturing the inherent temporal patterns in time-series data due to their reliance on the self-attention mechanism, and have high computational overhead.
 DLinear and NLinear models Zeng et al.
 (2023) take a simpler and more efficient approach that mitigates these issues by focusing on linear models that are more adept at handling the sequential nature of time-series data.
-One recent related work is CycleNet Lin et al.
-(2024), which introduces Residual Cycle Forecasting (RCF) to model periodic patterns using learnable recurrent cycles and predict residual components.
-However, CycleNet assumes stable periodicities, making it less effective for dynamic or irregular cycle datasets.
+To address the high computational overhead of transformer-based models, PatchTST Nie et al.
+(2022) segments time-series into patches, significantly improving computational efficiency and forecasting accuracy.
+Similarly, S-Mamba Wang et al.
+(2025) employs a selective state-space approach to efficiently capture inter-variate correlations and temporal dependencies with nearlinear complexity, achieving state-of-the-art performance across several benchmarks.
+Another recent work, CycleNet Lin et al.
+(2024), introduces Residual Cycle Forecasting (RCF) to model periodic patterns using learnable recurrent cycles and to predict residual components.
+However, CycleNet assumes stable periodicities, making it less effective for dynamic or irregular cycle datasets, i.e., datasets with regularly repeating patterns such as daily or seasonal patterns.
 Additionally, it models each time-series channel independently, failing to capture interdependencies between variables, which can limit its performance on multivariate time-series data.
-Graph Neural Networks (GNNs) have emerged as a powerful paradigm for time-series forecasting, particularly for multivariate time-series data.
-GNNs are designed to capture spatial dependencies by representing the data as graphs, where nodes represent variables, and edges represent dependencies between these variables.
-One notable approach is the MTGNN Wu et al.
-(2020) framework, which integrates temporal and graph convolution modules to capture both temporal and spatial dependencies, using a graph learning layer to adaptively construct an adjacency matrix that reflects hidden relationships among time-series variables.
-Another example is the Multivariate Time-Series Graph Ordinary Differential Equation (MTGODE) Jin et al.
-(2022) framework, which uses graph neural networks to model complex interactions among time-series and employs neural ODEs to continuously evolve the data over the graph structure, enhancing forecasting accuracy and robustness.
-Defining optimal graph structures and integrating graph learning with time-series forecasting in an end-to-end manner remains a significant challenge.
-While recent advances have introduced adaptive graph learning Zhang et al.
-(2020) and automated spatio-temporal fusion techniques Li and Zhu (2021), these methods come with their own limitations.
-Specifically, adaptive graph learning can be computationally expensive, and spatio-temporal fusion mechanisms may struggle to effectively capture long-range dependencies, leading to potential inaccuracies in forecasting Kumar et al.
-(2024).
-Recent advancements in time-series forecasting feature foundation models such as TimesFM Das et al.
+Recent advances in time-series forecasting feature foundation models such as TimesFM Das et al.
 (2024) and MOIRAI Woo et al.
@@
 However, these models can be computationally intensive, require significant resources for training and deployment, and often lack interpretability.
-Overall, while each category of time-series forecasting methods has distinct limitations, the integration of graph neural networks with deep learning techniques represents a promising direction to advance the state-of-the-art in multivariate time-series forecasting.
-3 Problem Formulation
-Let $X=(x_{1},x_{2},\ldots,x_{T})\in\mathbb{R}^{T\times d}$ represent a time-series of $T$ observations, each being $d$-dimensional.
+Graph Neural Networks (GNNs) have emerged as a powerful paradigm for multivariate time-series forecasting by modeling variables as nodes and their interdependencies as edges in a graph.
+The MTGNN Wu et al.
+(2020) framework introduces a joint architecture combining temporal and graph convolution modules, with a graph learning layer that adaptively constructs an adjacency matrix to uncover hidden variable relationships.
+In contrast, MTGODE Jin et al.
+(2022) employs neural ordinary differential equations over dynamically learned graphs to model continuous spatial-temporal dynamics, offering improved accuracy and robustness in forecasting tasks.
+Despite the large amount of work that utilizes GNNs for time-series forecasting, discovering optimal graph structures and integrating graph learning with time-series forecasting remains a significant challenge.
+While recent advances have introduced adaptive graph learning Zhang et al.
+(2020) and automated spatio-temporal fusion techniques Li and Zhu (2021), these methods come with their own limitations.
+Specifically, adaptive graph learning can be computationally expensive, and spatio-temporal fusion mechanisms may struggle to effectively capture long-range dependencies, leading to potential inaccuracies in forecasting Kumar et al.
+(2024).
+To evaluate the proposed GNN-based long-term forecasting model, we consider the best model from each of the above categories—namely linear models, state-space models, transformer-based models, and foundation models—as our baseline.
+3 Methodology
+Let $X=(x^{1},x^{2},\ldots,x^{T})\in\mathbb{R}^{T\times D}$ represent a time-series of $T$ observations, each being $D$-dimensional, i.e.
+emitted by $D$ sensors.
 Time-series forecasting concerns predicting future values based on past observations.
-Given a look-back window of length $l$ ending at time $t$, denoted as $X_{t,l}=(x_{t-l+1},\ldots,x_{t-1},x_{t})$, the goal is to forecast the next $h$ steps of the time-series.
-This is achieved by learning the function $f_{\theta}(X_{t,l})=(\hat{x}_{t+1},\ldots,\hat{x}_{t+h})$, where $\theta$ represents the parameters of the forecasting model.
+Given a look-back window of length $l$ ending at time $t$, denoted as $X^{t,l}=(x^{t-l+1},\ldots,x^{t-1},x^{t})$, the goal is to forecast the next $h$ steps of the time-series.
+This is achieved by learning the function $f_{\theta}(X^{t,l})=Y^{t,h}=(y^{t+1},\ldots,y^{t+h})$, where $\theta$ represents the parameters of the forecasting model.
 We refer to each pair of look-back and forecast windows as a sample.
-4 Methodology
-We model the input and output time-series as a graph.
-Each sensor channel $i$ is represented as a node $v_{i}$ that might have relationships with other channels, represented as directed and weighted edges $e_{ij}$ connecting these nodes.
-These connections are assumed to be static, i.e., the graph structure $G=(V,E)$ does not change over time.
-The node features are initially the historical values of the corresponding channels $x_{i}=\{x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{T}\}$, and ultimately the forecast values $\hat{y}_{i}=\{\hat{y}_{i}^{T+1},\hat{y}_{i}^{T+2},\ldots,\hat{y}_{i}^{T+H}\}$.
-The proposed methodology has two main parts:
+Since we focus on long-term forecasting, the forecast window is assumed to be long, usually several days to a few months.
+We model the input and output time-series as a graph $G=(V,E)$.
+Each sensor $i$ is represented as a node $v_{i}$ that might have dependencies with other sensors.
+These dependencies are modeled using directed, weighted edges connecting the respective nodes.
+These connections are assumed to be static, i.e., the graph structure does not change over time.
+The node features are initially the historical sensor readings with the look-back window, and ultimately the forecast values.
+Our methdology has three main steps:
   1. 1.
-    Graph Structure Learning: This involves learning the adjacency matrix $A$ capturing the relationships between nodes.
-The matrix $A\in\mathbb{R}^{D\times D}$ where $D$ is the number of nodes, and each element $a_{ij}$ represents the weight of the edge from node $v_{i}$ to node $v_{j}$.
+    Decomposition: Decomposing the time-series into fundamental components (trend, seasonal, residual) to accurately capture inherent characteristics of each component.
   2. 2.
-    Spatio-Temporal Forecasting: This is the core of our model, responsible for learning and prediction. It comprises three subcomponents:
-    1. (a)
-      Decomposition: Each input sample is decomposed into its fundamental components to better capture underlying patterns.
-    2. (b)
-      Spatio-Temporal Feature Extraction: This step extracts features from the graph using both spatial and temporal information.
-For each node, $v_{i}$, the feature vector $h_{i}$ is derived by combining its historical values $x_{i}$ and the influence of its neighboring nodes as defined by $A$.
-    3. (c)
-      Forecasting: Finally, the model uses the extracted features to predict future values.
-The forecast values $\hat{Y}$ for all nodes are given by $\hat{Y}=f(H,A)$, where $H$ is the matrix of all node features and $f$ is the forecasting function.
-Each of these parts will be explained in more detail in the subsequent sections.
+    Graph Structure Learning: Learning the adjacency matrix $A\in\mathbb{R}^{D\times D}$ (with $D$ being the number of nodes) that captures the relationships between nodes for each time-series component.
+Each element of $A$, denoted as $a_{ij}$, represents the weight of the edge from node $v_{i}$ to node $v_{j}$.
+  3. 3.
+    Spatio-Temporal Forecasting via DST: Utilizing the learned graph structure for each component, a GNN extracts spatio-temporal features for predicting future values of the respective component.
+The predicted future values of different components are then summed to produce the final forecast $Y^{t,h}$.
+Each of these steps will be explained in more detail in the subsequent sections.
+3.1 Decomposition
+The first step in our approach involves decomposing the time-series into its fundamental components: trend, residual, and one or more seasonal components.
+This decomposition is crucial as it isolates the inherent characteristics of the time-series, thus improving the accuracy of forecasting Sohrabbeig et al.
+(2023).
+The decomposition method adopted here is inspired by the MSTL decomposition technique Bandara et al.
+(2025).
+It begins with the application of a weighted moving average filter to the time-series, where the weights for this kernel are derived from the tricube function given below:
+$\text{tricube}(x)=\begin{cases}(1-|x|^{3})^{3},&\text{if }|x|\leq 1,\\ 0,&\text{if }|x|>1.\end{cases}$ (1)
+Intuitively, it assigns higher weights to nearby data points and lower weights to distant ones. The extracted trend is then subtracted from the original time-series to obtain the detrended series.
+Next, the period of seasonality (e.g., daily or weekly patterns) must be determined.
+This can be done using various methods, such as autocorrelation analysis, Fourier transform, data visualization, etc.
+The next step is to capture seasonal patterns in the detrended series.
+A simple aggregation approach with a sliding window of appropriate length (denoted as $k$) and stride (denoted as $m$) is used for this purpose.
+The seasonal pattern is identified by averaging the segments covered by the sliding window.
+Figure 1: Extracting seasonal patterns using a sliding window of size $k$ and a stride of $m$ times the length of the seasonality period, respectively. In this example, $k=3$ and $m=1$.
+Figure 1 illustrates this process. In each step of moving the sliding window, $S_{i}$ is calculated using the following equation:
+$S_{i}=\frac{1}{k}\sum_{j=i-\lfloor k/2\rfloor}^{i+\lfloor k/2\rfloor}B_{i},$ (2)
+where $B_{i}$ is the i-th segment of the detrended time-series.
+Each segment contains $p$ data points, where $p$ represents the length of the seasonality period inferred in the previous step.
+Hence, the data points in $B_{i}$ have indices $i\times p$ to $(i+1)\times p-1$.
+To preserve the input length, padding is added to both ends of the time-series.
+This padding consists of copies of the first and last blocks.
+Finally, the so-obtained seasonal components are subtracted from the detrended series to obtain the residual component.
 3.2 Graph Structure Learning
-Since the true underlying graph is unknown, we infer it from the data.
-To this end, we first decompose the time-series into its fundamental components and then infer the underlying graph for each component independently.
-This entire process is conducted offline.
-The graph structure is assumed to be static, hence it is not updated over time.
-4.1.1 Decomposition
-The first step in our approach involves decomposing the time-series into three fundamental components: trend, residual, and one or more seasonal components.
-This decomposition is crucial as it isolates the inherent characteristics of the time-series, thus improving the accuracy of forecasting and analysis Sohrabbeig et al.
-(2023).
-The decomposition process begins with the application of a moving weighted average filter to the time-series.
-The weights for this kernel are derived from the tricube function given below:
-$\text{tricube}(x)=\begin{cases}(1-|x|^{3})^{3},&\text{if }|x|\leq 1,\\ 0,&\text{if }|x|>1.\end{cases}$ (1)
-It assigns higher weights to nearby data points and lower weights to distant ones. The extracted trend is then subtracted from the original time-series to obtain the detrended series.
-Next, the period of seasonality (e.g., daily or weekly patterns) must be determined.
-The period of seasonality can be identified using methods such as autocorrelation analysis, Fourier Transform, data visualization, or expert prior knowledge; in this work, we relied on data visualization and prior expert knowledge.
-To capture seasonal patterns in the detrended series, an aggregation approach with a sliding window of appropriate length and stride is used.
-The seasonal pattern is identified by averaging the segments covered by the sliding window.
-Fig.
-1 illustrates this process.
-Figure 1: Extracting seasonal patterns using a sliding window of size $l$ and a stride of $m$ times the length of the seasonality period, respectively. In this example, $l{=}3$ and $m{=}1$.
-In each step of moving the sliding window, $S_{i}$ is calculated using the following equation:
-$S_{i}=\frac{1}{l}\sum_{j=i}^{i+l-1}B_{i},$ (2)
-where $B_{i}$ is the i-th block of the detrended time-series.
-Each block is a vector containing $p$ data points, where $p$ represents the length of the seasonality period, ranging from index $i\times p$ to $(i+1)\times p-1$.
-To preserve the input length, padding is added to both ends of the time-series.
-This padding consists of repeated copies of the first and last blocks.
-Finally, the so-obtained seasonal components are subtracted from the detrended series to obtain the residual component.
-Each variate of the multivariate time-series, except for date and time features, undergoes this decomposition process separately.
-By the end of this procedure, we have extracted a trend, a residual, and one or multiple seasonal components for each variate, ensuring that the unique characteristics of each time-series are accurately captured and ready for the subsequent analysis.
-4.1.2 Learning graph structure for each component
-After decomposing the time-series into its constructing components, we construct the underlying graph for each component by connecting a subset of nodes.
-To ensure computational efficiency, each component is downsampled to a level that preserves its general patterns.
-Next, the distance between two variates is measured using the Dynamic Time Warping (DTW) Müller (2007) metric.
-DTW is chosen for its robustness in handling nonlinear alignments, providing reliable similarity measurements between time-series.
-Finally, we connect each node to its $K$ nearest neighbors according to the DTW distance, with $K$ being a hyperparameter in our approach.
-The described procedure is outlined in Alg.
-1.
-Algorithm 1 Time-series Graph Construction
-0:  time-series data $\{X_{i}\}_{i=1}^{D}$, seasonality periods $P_{1},P_{2},...,P_{r}$, number of nearest neighbors K
-0:  Underlying Graph G
-1:  Pattern Extraction:
-2:  for each time-series $X_{i}$ do
-3:     Extract trend component $T_{i}$ using a moving weighted average with tricube weights and then downsample
-4:     Detrend the series: $\hat{X}_{i}=X_{i}-T_{i}$
-5:     Extract seasonality component $S_{i}$
-6:     for each seasonality period $P$ in $\{P_{1},P_{2},...,P_{r}\}$ do
-7:        Apply a moving aggregation approach with window length $l\times P$ and stride $m\times P$
-8:        Compute seasonality component $S_{i}$ as the average of the $n$ periods
-9:     end for
-10:  end for
-11:  Distance Metric:
-12:  Use DTW metric to calculate the distance between any two pairs of variates and construct distance matrix $D$ as shown in Fig. 3
-13:  Construct the Underlying Graph
-14:  Connect each node to its K nearest neighbors
-Fig.
-2 depicts the inferred graph from the trend component of the carbon intensity dataset, described in Section 5.1.
-Here the inferred graph is directional, as one variate might be dependent on the lagged version of the other one, but the opposite is not necessarily true.
-Fig.
-3 illustrates the Dynamic Time Warping (DTW) distances between different nodes for both the trend and seasonal components of the air pollution dataset.
-These heatmaps reveal distinct patterns of similarity, with lower normalized DTW values indicating stronger temporal alignment between nodes.
-Figure 2: The carbon intensity value is shown for a subset of states in the US at a given point in time.
-Dark brown (light yellow) indicates higher (lower) carbon intensity in a jurisdiction.
-Data has been collected from a number of power generation companies, which may supply power to different states.
-Consequently, the relationship between the states and energy-providing companies is many-to-many.
-Some arrows extend beyond the map because the image shows only a portion of the full area, and some zones connect to others not included in the display.
-Figure 3: Heatmap of DTW distances between different nodes for (a) trend and (b) seasonal components of the air pollution dataset, described in Section 5.1.
-Note that the distance matrices are symmetrical, with lower values indicating greater closeness.
-All values are min-max normalized to ensure consistent scaling.
-4.2 Spatio-Temporal Forecasting
+Since the underlying graph structure for each time-series component is not known in advance, we have to infer it from historical data.
+To this end, we first apply the decomposition process described in the previous section to the entire training dataset, then independently construct a graph for each resulting component.
+We note that graph structure learning is an offline task as the graph structures are assumed to be static.
+The overall procedure is summarized in Algorithm 1.
+It starts with decomposing the time-series into its fundamental components, where the number of these components is assumed to be identical across all variates.
+To ensure computational efficiency, each component is first downsampled to a level that preserves its essential structural patterns.
+To quantify similarity between sensor readings within each time-series component, we compute pairwise distances between the variates using Dynamic Time Warping (DTW) Müller (2007).
+DTW is selected for its robustness in capturing nonlinear temporal alignments between time-series.
+Each node is then connected to its $K$ nearest neighbors according to the DTW distances, where $K$ is a hyperparameter.
+Algorithm 1 Graph Structure Learning for Time-Series Components
+0: time-series data $\{X_{i}\}_{i=1}^{D}$, number of nearest neighbors $K$
+0: graph $G_{i}$ for each time-series component $i$
+1:for each time-series variate $X_{i}$ do
+2:  $\hat{X}_{i}\leftarrow\text{Detrend}(X_{i})$
+3:  $\{P_{1},P_{2},...,P_{n}\}\leftarrow\text{GetSeasnonalityPeriods}(\hat{X}_{i})$
+4:  for each seasonality period $P_{j}$ in $\{P_{1},P_{2},...,P_{n}\}$ do
+5:   Compute seasonal component $S_{j}$
+6:  end for
+7:  Compute the residual component
+8:end for
+9:for each extracted component of $X$ denoted as $T_{i}$ do
+10:  Downsample $T_{i}$
+11:  Build matrix $A_{i}$ by computing DTW between variates
+12:  Construct $G_{i}$ by selecting top $K$ values in $A_{i}$
+13:end for
+Figures 2(a) and 2(b) illustrate DTW distances between different nodes for both the trend and seasonal components of the air pollution dataset, described in Section 4.1.
+These heatmaps reveal distinct patterns of similarity, with lower normalized DTW values indicating stronger dependencies between nodes.
+Thus, the graphs constructed for the trend and seasonal components have different structures.
+Our ablation study, described in Section 6, confirms that this results in a significant improvement over a GNN-based forecasting model that does not perform decomposition or uses the same graph structure for all components.
+Figure 2(c) depicts the graph constructed for the trend component of the carbon intensity dataset, described in Section 4.1.
+As it can be seen, the inferred graph contains directed edges, as one variate might be dependent on the lagged version of the other one, but the opposite is not necessarily true.
+Figure 2: DTW distance heatmaps of (a) trend and (b) seasonal components of the air pollution dataset (Section 4.1); lower values indicate greater similarity, with all values min-max normalized.
+(c) Carbon intensity for selected US states at a given time, where darker colors indicate higher intensity.
+Data come from multiple power generation companies with many-to-many links to states; some arrows extend beyond the displayed map to conneced areas not shown.
+3.3 Spatio-Temporal Forecasting via DST
 Figure 3: The overall architecture of proposed forecasting model. After decomposition, each component is handled independently.
-Now that we have learned the respective graph structure for each constituent component in an offline fashion, as described in Section 4.1.1, we proceed with the spatio-temporal forecasting step.
-The overall architecture of the proposed forecasting model is illustrated in Fig.
-4.
-The input of the model is a sample of fixed-length historical values.
-These historical values include date and time features, along with a number of variates that we would like to predict.
+After constructing a graph for each time-series component in an offline fashion, we proceed with spatio-temporal forecasting.
+The architecture of the proposed forecasting model, DST, is shown in Figure 3.
+The model’s input is a sample of fixed-length historical values, $X^{t,l}$, and the corresponding date-time values, denoted as $X_{\text{time}}$.
 3.3.1 Sample Decomposition
-Figure 5: Decomposition of an electricity consumption sample over two weeks.
-The first subplot displays the original data alongside the sum of the trend and seasonal components, showing their closeness.
-First, we need to decompose an input sample, without date and time features, into its constituent components: trend, seasonality, and residual.
-This process follows the procedure described in Section 4.1.1, but without the need for downsampling.
-Fig.
-5 illustrates the decomposition of a randomly selected sample from the Electricity dataset, described in Section 5.1.
-4.2.2 Spatial feature extraction
-Our feature extraction model handles each component, that is, trend, seasonal, and residual, independently.
-For each component, we extract and combine three types of features, namely date-time, spatial, and temporal features.
-The date-time features are processed through a linear layer to capture some temporal patterns.
-Simultaneously, spatial and temporal features are extracted from the input data.
-These temporal and spatial features are then combined and passed on to the forecasting module that uses them to generate the final output.
-Fig.
-6 is a schematic diagram of the feature extraction module.
-$\hat{y}_{t}=f_{\text{forecast}}\left(f_{\text{time}}(x_{\text{time}})\oplus f_{\text{spatio-temporal}}(x)\right)$ (3)
+Figure 4: Decomposition of an electricity consumption sample over two weeks. The first subplot displays the original data alongside the sum of trend and seasonal components.
+We begin by decomposing an input sample, without the date-time series, into its constituent components: trend, seasonality, and residual.
+For decomposition, we follow the same procedure described in Section 3.1, but without downsampling.
+A separate forecasting model is trained for each component using the graph structure learned for that component specifically.
+Figure 4 illustrates the decomposition of a randomly selected sample from the Electricity dataset, described in Section 4.1.
+In this case, there is just one seasonal component.
+3.3.2 Feature extraction
+The first module of DST is responsible for feature extraction.
+This module handles each component, i.e., trend, seasonal, and residual, separately.
+Specifically, for each component, we extract and then combine three types of features, namely spatial and temporal features and date-time embedding.
+The result is passed as input to the forecasting module, which generates the final output for that component.
+Figure 5 shows DST’s feature extraction and forecasting modules.
+To make the date-time embedding, we form a series ($X_{\text{time}}$) by putting the timestamps of input data points in a sequence.
+We then pass this to an encoder that uses 1D convolution to produce the date-time embedding.
+Simultaneously, spatial and temporal features are extracted from the input data $X^{t,l}$.
+These temporal and spatial features are then combined and fed to the forecasting module as written below:
+$Y^{t,h}=f_{\text{forecast}}\left(f_{\text{time}}(X_{\text{time}})\parallel f_{\text{spatio-temporal}}(X^{t,l})\right),$ (3)
+where $\parallel$ is the concatenation operator, $f_{\text{forecast}}$ is the forecasting model, $f_{\text{time}}$ is the date-time encoder, and $f_{\text{spatio-temporal}}$ is the spatiotemporal feature extractor, which is a combination of GATv2 and TCN as described next.
 Figure 5: One layer of spatio-temporal feature extraction with time embedding.
-The input to this module has n channels, $C_{0}$ to $C_{n}-1$.
-Each component from the decomposition module—trend, seasonal, and residual—is processed independently using this layer.
-The spatial feature extractor is a Graph Attention Network v2 (GATv2), with a pre-inferred, static underlying graph structure that does not change over time.
-Figure 7: Modeling of multivariate time-series as a graph. Each node represents a variate, and its historical data serves as the features of that node
-For spatial feature extraction, we utilize a model based on GATv2 Brody et al.
-(2021).
-In this model, each variate of the time-series is treated as a node, with its historical values within the sample serving as its features, as depicted in Fig.
-7.
-GATv2 requires a connection graph to understand the relationships between different nodes (variates).
+The input to this module is readings from $n$ sensors, $X_{0}$ to $X_{n-1}$.
+Each component from the decomposition module is processed separately using this layer.
+The spatial feature extractor is a GATv2 that relies on the learned graph structure.
+For spatial feature extraction, we utilize a GNN.
+In this model, each variate of the time-series is treated as a node, with its historical values within the look-back window serving as its features.
+The GNN requires a connection graph to understand the relationships between different nodes (variates) via message passing.
 The connection graph is inferred using the method described in Section 3.2.
-In a GNN, each node iteratively updates its state by aggregating information from its neighbors. The general formulation of a GNN layer is as follows:
-$\mathbf{h}_{i}^{\prime}=\sigma\left(\mathbf{W}\mathbf{h}_{i}+\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j}\mathbf{h}_{j}\right)$ (4)
-where $\mathbf{h}_{i}$ is the feature vector of node $i$, $\mathcal{N}_{i}$ represents the set of neighbors of node $i$, $\mathbf{W}$ and $\mathbf{W}_{j}$ are learnable weight matrices, and $\sigma$ is a nonlinear activation function.
-This aggregation allows the model to capture the dependencies between nodes and their neighbors, which is crucial for spatial feature extraction.
-In GNNs, Equation 4 is used to calculate the updated node representation as a weighted sum of the neighbors’ representation.
-To increase learning capacity, we can incorporate this method with an attention mechanism.
-The Graph Attention Network (GAT) Veličković et al.
-(2017) enhances the GNN framework by incorporating an attention mechanism.
-This mechanism allows nodes to assign different amounts of importance to the neighbors, which enables capturing of changing dynamics.
-Specifically, the attention score $e_{ij}$ between node $i$ and its neighbor $j$ is computed as:
-$e_{ij}=\text{LeakyReLU}\left(\mathbf{a}^{T}[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}]\right)$ (5)
+In a GNN, each node iteratively updates its state by aggregating information from its neighbors. The general formulation of a GNN layer is given below
+$\mathbf{h}_{i}^{\prime}=\sigma\left(\mathbf{W}\mathbf{h}_{i}+\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j}\mathbf{h}_{j}\right),$ (4)
+where $\mathbf{h}_{i}$ is the feature vector of node $i$, $\mathcal{N}_{i}$ represents the set of neighbors of node $i$ in the underlying graph, $\mathbf{W}$ and $\mathbf{W}_{j}$ are learnable weight matrices, and $\sigma$ is a nonlinear activation function.
+This aggregation allows the model to capture the dependencies between each node and its neighbors, which is crucial for spatial feature extraction.
+The output $\mathbf{h}_{i}^{\prime}$ is the updated representation of node $i$, which is a weighted sum of its neighbors’ representations.
+To increase learning capacity, we use the attention mechanism for aggregation.
+Specifically, we use GATv2 Brody et al.
+(2021), a dynamic graph attention variant of GAT, to allow nodes to assign different amounts of importance to their neighbors in the aggregation step.
+This enables capturing time-varying dependecies.
+The attention score $e_{ij}$ between node $i$ and its neighbor $j$ is computed as:
+$e_{ij}=\mathbf{a}^{T}\text{LeakyReLU}\left(\mathbf{W}[\mathbf{h}_{i}\parallel\mathbf{h}_{j}]\right)$ (5)
 where $\mathbf{a}$ is a learnable weight vector, $\mathbf{W}$ is a weight matrix, $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the feature vectors of nodes $i$ and $j$, respectively, and $\parallel$ denotes concatenation.
@@
 $\alpha_{ij}=\text{softmax}_{j}(e_{ij})=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$ (6)
-The final output representation of node $i$ is computed as a weighted sum of its neighbors’ features:
+The final output representation of node $i$ is computed by passing a weighted sum of its neighbors’ features to the activation function:
 $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (7)
-This attention mechanism allows GAT to focus on the most relevant neighbors, improving its ability to capture complex relationships in the graph.
-GATv2 Brody et al.
-(2021), builds on the original GAT by addressing its limitations and introducing a more expressive dynamic attention mechanism.
-In GATv2, the attention score computation is modified to condition both the query and key nodes, enhancing the model’s flexibility.
-The new attention score $e_{ij}$ is defined as:
-$e_{ij}=\mathbf{a}^{T}\text{LeakyReLU}\left(\mathbf{W}[\mathbf{h}_{i}\parallel\mathbf{h}_{j}]\right)$ (8)
-In this architecture, we employed just one layer of GATv2.
-The decision of whether to use one layer of GATv2 or stack more of them has been made empirically.
-Fig.
-9 in Appendix A justifies why one layer of the GATv2 layer was employed for the carbon intensity dataset.
-4.2.3 Temporal Feature Extraction
+This attention mechanism allows the GNN to focus on the most relevant neighbors, improving its ability to capture complex relationships in the graph.
+We use just one GATv2 layer for spatial feature extraction.
 For temporal feature extraction, we utilize Temporal Convolutional Networks (TCNs) Bai et al.
-(2018).
-TCNs are designed to capture temporal dependencies in sequential data by leveraging convolution across time steps.
-The key components of TCNs are causal convolutions, dilations, and residual connections.
-Causal convolutions ensure that the model only uses past information to predict future values, preserving the temporal order of the data. The causal convolution for a given layer $l$ is defined as:
-$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i}$ (9)
-where $\mathbf{y}_{t}^{l}$ is the output at time $t$ for layer $l$, $\mathbf{W}_{i}^{l}$ are the convolutional filters, $\mathbf{x}_{t-i}$ are the inputs from the previous layer, and $k$ is the filter size.
-Dilated convolutions expand the receptive field of the network without increasing the number of parameters, allowing the model to capture long-range dependencies.
-The dilation factor $d$ determines the spacing between the kernel elements.
-A dilated convolution is defined as:
-$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i\cdot d}$ (10)
-where $d$ is the dilation factor, $k$ is the filter size, and the other symbols retain their previous meanings.
-Residual connections help mitigate the vanishing gradient problem and facilitate the training of deeper networks. The output of each TCN layer is combined with its input through a residual connection:
-$\mathbf{y}_{t}=\sigma\left(\mathbf{y}_{t}^{l}+\mathbf{x}_{t}\right)$ (11)
+(2018), which have demonstrated strong performance on sequential data across multiple domains.
+Our choice of TCN is motivated both by empirical results and by its architectural advantages over alternatives like LSTM and GRU.
+TCNs capture temporal dependencies using convolutional operations over time steps, which offer parallelism during training, stable gradients, and a flexible receptive field.
+Compared to RNN-based models, such as LSTM and GRU, which process data sequentially and may struggle with long-range dependencies, TCNs enable faster training and are better suited to modeling long sequences, a key challenge in long-term time-series forecasting.
+The adopted TCN has three core components Bai et al.
+(2018): causal convolutions, dilated convolutions, and residual connections.
+Causal convolutions ensure that predictions at time $t$ depend only on past inputs, preserving the temporal order.
+Dilated convolutions expand the receptive field without increasing model size, enabling the network to capture long-range patterns.
+Residual connections promote efficient training of deeper models and help mitigate vanishing gradients.
+Overall, TCNs offer a strong balance between expressiveness, training efficiency, and robustness for long-range temporal modeling, making them especially well-suited to the demands of multivariate, long-term forecasting of urban data.
 3.3.3 Forecasting
-Once features are extracted for each component of the time-series, we employ a linear forecasting model to produce forecasts for the entire forecast horizon at once, utilizing a direct multistep forecasting approach.
+Once features are extracted for each component of the time-series, we employ a linear forecasting model to produce forecasts for the entire forecast horizon at once; this approach is commonly referred to as direct multistep forecasting.
 This choice is motivated by the demonstrated effectiveness of similar methods in previous studies Sohrabbeig et al.
@@
 Forecasts are generated simultaneously and independently for each component—trend, seasonal, and residual—mirroring the approach used in the feature extraction phase.
-Within each forecasting component, each variate is forecasted in a univariate fashion.
+Within each forecasting component, each variate is predicted in a univariate fashion.
 This means that only the spatial and temporal features extracted for each specific variate in the previous step are used for forecasting that variate.
-Nevertheless, these features are influenced by the features of some other variates due to the way that they are generated.
-The input to the forecasting module is the combination of spatio-temporal features and time embeddings.
-In this work, these features and embeddings are added together before being fed into the forecasting task.
-The decision to add rather than concatenate the time embeddings with the other extracted features was made based on our experimental results.
-Fig.
-6 shows the input for the forecasting module.
-The linear forecasting model operates on input data with the following characteristics:
-  • Input Sequence Length ($L$): The length of the input time-series sequence.
-  • Prediction Length ($H$): The number of future time steps to be predicted.
-  • Number of Channels ($D$): The number of separate time-series channels in the input data.
+Nevertheless, these features are influenced by the features of other variates due to the way that they are extracted using a GNN.
+The input to the forecasting model is the combination of spatio-temporal features and the time embedding (see Equation 3).
+Figure 5 shows how the input for the forecasting model is constructed.
+We adopt a linear forecasting model that predicts values in the next $h$ time steps.
 The input to the model is a tensor $\mathbf{X}$ of shape $(B,D,l)$, where $B$ is the batch size, $D$ is the number of channels, and $l$ is the length of the input sequence.
 During the forward pass, the model computes the prediction $\mathbf{Y}$ using the input tensor $\mathbf{X}$ and the learnable parameters.
-This is done through a linear transformation followed by the addition of the bias term:
-$\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}+\mathbf{b}$ (12)
+This is done through a linear transformation followed by the addition of a bias term:
+$\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}+\mathbf{b},$ (8)
 where $\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}$ represents the Einstein summation operation defined as:
@@
 (2023) models, which iteratively calculates forecasts for each variate.
-5 Result
-In this section, we first introduce the datasets and baseline models used for evaluation, then evaluate and compare the performance of our proposed model with the baselines in three different settings.
+3.3.4 Training
+As explained earlier, DST takes advantage of a dedicated module for each component that was extracted via the decomposition method.
+While each module processes the respective component independently, all modules are jointly trained.
+For training, we used the PyTorch Lightning framework on a high-performance computing system featuring an AMD Ryzen 7 5800X 8-core processor (16 threads), 32GB RAM, and an NVIDIA GeForce RTX 3090 GPU with 24GB VRAM.
+The model is optimized using the Adam optimizer with a learning rate of 0.0001 and a batch size of 32.
+Training is performed for a maximum of 20 epochs, with early stopping applied based on validation loss, using a patience of 3 epochs to mitigate overfitting.
+Gradient clipping with a threshold of 1.0 is employed to ensure training stability.
+The Mean Squared Error (MSE) loss function is used for training DST, while both MSE and Mean Absolute Error (MAE) are reported during evaluation to provide a comprehensive assessment of forecasting performance.
+We note that the hyperparameter K, which defines the number of nearest neighbours for each node in the graph, is selected using grid search over a range spanning from 10% to 100% of all other nodes, based on validation performance.
+4 Evaluation
+We introduce the urban datasets and baseline models used for evaluating the long-term forecasting performance of DST.
 4.1 Datasets
-For comprehensive coverage of critical factors influencing time series data, we select four urban datasets including electricity usage, weather conditions, carbon intensity, and air pollution.
-These datasets provide a robust benchmark for understanding the pros and cons of different time series forecasting models.
-More details on these datasets are provided below:
-Table 1 summarizes all the main characteristics of the above datasets.
-With the exception of the air pollution dataset, the other datasets did not require any imputation of missing values.
-For the air pollution dataset, missing values were filled using the average of records corresponding to the same month, day of the week, and hour.
-As a preprocessing step, all columns in all datasets were standardized.
-Table 1: Statistical details of the real-world datasets used in this study.
-Dataset No. Features Length Time Resolution
+To comprehensively evaluate our model in realistic and diverse long-term forecasting scenarios, we utilize four multivariate time-series datasets spanning key domains: electricity usage, weather conditions, carbon intensity, and air pollution.
+These datasets were carefully selected to reflect a wide range of forecasting challenges, such as high dimensionality, varying temporal resolutions, and complex inter-variable dependencies.
+As a result, they are suitable benchmarks for assessing the generalizability and robustness of multivariate time-series forecasting models.
+We describe each dataset below:
+  • Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014: This dataset consists of hourly electricity consumption data from 321 customers over a period of three years.
+It serves as a benchmark for modeling periodic, high-dimensional energy demand, where capturing temporal patterns and seasonal cycles is critical.
+  • Weather [^2] https://www.bgc-jena.mpg.de/wetter/: This dataset includes 21 meteorological metrics collected at 10-minute intervals throughout 2020 from a weather station at the Max Planck Biogeochemistry Institute.
+Its high-frequency nature makes it ideal for evaluating models’ sensitivity to rapid fluctuations and short-term dependencies.
+  • Carbon Intensity [^3] https://www.electricitymaps.com/data-portal/united-states-of-america: This dataset contains hourly carbon intensity values recorded across 50 U.S.
+states and four territories over three years, this dataset introduces spatio-temporal diversity and is crucial for applications involving environmental sustainability and energy emissions forecasting.
+  • Air Pollution [^4] https://archive.ics.uci.edu/dataset/381/beijing+pm2+5+data: This dataset provides hourly PM2.5 concentrations recorded by the U.S.
+Embassy in Beijing, along with concurrent meteorological data from the Beijing Capital International Airport Liang et al.
+(2015).
+It presents a complex, multivariate forecasting task influenced by both environmental and anthropogenic factors.
+These datasets collectively span different temporal granularities, input complexities, and domain characteristics, making them an effective testbed for evaluating models intended for urban forecasting.
+Moreover, they align with benchmarks used in recent foundational and graph-based forecasting research Zeng et al.
+(2023); Wang et al.
+(2025), ensuring relevance and comparability.
+Table 1 summarizes the key characteristics of each dataset.
+There are no missing values in these datasets, except for the air pollution dataset, which required imputation using averages of similar temporal contexts (same month, weekday, and hour).
+As a preprocessing step, we standardized all features.
+Dataset # Variates Length Time Resolution
 Electricity 321 26,304 1h
 Air Pollution 7 34,970 1h
-Carbon 54 26,280 1h
+Carbon Intensity 49 26,280 1h
 Weather 21 52,696 10min
-5.2 Baselines
-To evaluate our model, we use three highly effective Transformer models, namely Informer Zhou et al.
-(2021), Autoformer Wu et al.
-(2021), and FEDformer Zhou et al.
-(2022), as our baselines.
-Additionally, we consider two recent state-of-the-art models, NLinear and DLinear Zeng et al.
+Table 1: Statistical details of the real-world datasets used in this study.
+4.2 Baseline Forecasting Models
+Following the literature review provided in Section 2, we select representative state-of-the-art baselines across multiple forecasting paradigms for comprehensive evaluation of our approach.
+Transformer-based models include Autoformer Wu et al.
+(2021) and PatchTST Nie et al.
+(2022), both of which are designed to effectively capture temporal patterns through attention mechanisms.
+Autoformer is chosen over early adaptations of transformer-based models for time-series forecasting, namely Informer Zhou et al.
+(2021) and Fedformer Zhou et al.
+(2022), primarily due to its consistently stronger performance on our datasets, making it a more suitable representative of that class of models.
+From the class of linear models, we incorporate DLinear Zeng et al. (2023), known for its efficiency and competitive performance on long-sequence forecasting tasks.
+To assess the capabilities of large-scale pretrained models, we consider TimesFM Das et al.
+(2024), evaluated in a zero-shot inference setting without fine-tuning, thereby highlighting the model’s generalization ability.
+MOIRAI Woo et al.
+(2024) is excluded from our evaluation as it is primarily designed for probabilistic forecasting, whereas our setup focuses on point forecasting.
+Additionally, we evaluate against Simple-Mamba (S-Mamba) Wang et al.
+(2025), a recent selective state-space model that leverages a lightweight architecture to efficiently capture both inter-variate correlations and temporal dependencies with near-linear computational complexity.
+We also include Repeat Last, a naive baseline that simply repeats the last observed value. This serves as a reference for assessing the forecasting performance of other models.
+Classical statistical models such as ARIMA and seasonal ARIMA (SARIMA) are excluded from direct comparisons due to their high computational cost for parameter selection and their consistently poor performance on long sequence time-series forecasting tasks, as reported in prior studies Wu et al.
+(2021); Zhou et al.
 (2021).
-We also employed a straightforward baseline: Repeat Last, which replicates the last observation.
-Our next baseline is TimesFM Das et al.
-(2023), a foundation model for time series forecasting that leverages a decoder-only attention mechanism with input patching.
-TimesFM is pre-trained on a large corpus of both real-world and synthetic time series data, enabling it to achieve near state-of-the-art zero-shot performance across various datasets without the need for fine-tuning.
-In these experiments, TimesFM was used solely for inference and was not fine-tuned.
-In this study, we did not compare our results with classical models such as ARIMA and SARIMA, due to the extensive time required for their parameter selection, which significantly slows down the training process.
-Instead, we based our conclusions on experimental results reported in other studies Wu et al.
-(2021); Zhou et al.
-(2021), which have shown that these classical models perform poorly on the long sequence time series forecasting (LSTF) task.
-Therefore, we opted not to include them in our direct comparisons.
-5.3 Experiments
-We now evaluate the performance of our model (labelled DST in the tables) and compare it with strong baselines in three different settings.
+Moreover, we do not consider models that have been specifically designed for a particular domain or dataset (e.g.
+air pollution or carbon intensity) as part of our baseline comparisons.
+This is because they cannot be readily used in other domains, often requiring changes to their architecture or pre-processing steps.
+5 Results
+We now evaluate the performance of our model (labelled DST in Table 2) and compare it with strong baselines in three different settings.
 In all experiments, the length of the look-back window is 336, and the learning rate is set to 0.0001.
-The length of the forecast horizon is denoted in the second column of tables 2, 3 and 4.
+The length of the forecast horizon is specified in the second column of Table 2.
 For our model, we decided to extract only one seasonal component due to the short length of the input.
 For instance, when the length of the look-back window is 336 hours (equivalent to two weeks), it is logical to extract daily seasonal patterns since multiple daily cycles are present in this period while extracting weekly patterns would be less meaningful considering the limited number of weekly cycles.
-Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.402 0.322 0.421 0.324 0.441 0.334 0.481 0.382 0.532 0.429 0.745 0.589 4.51 0.62 192 0.422 0.333 0.439 0.340 0.447 0.344 0.483 0.376 0.535 0.431 0.813 0.615 3.87 2.06 336 0.427 0.341 0.447 0.350 0.467 0.358 0.488 0.382 0.542 0.431 0.804 0.623 4.47 2.57 720 0.423 0.347 0.469 0.375 0.472 0.379 0.452 0.342 0.549 0.432 0.827 0.644 9.80 7.46 Carbon 96 0.468 0.501 0.476 0.503 0.467 0.492 0.512 0.539 0.519 0.547 1.128 0.780 -0.21 -1.83 192 0.534 0.545 0.541 0.544 0.540 0.547 0.674 0.621 0.584 0.580 1.243 0.829 1.11 0.37 336 0.589 0.581 0.601 0.595 0.597 0.594 0.703 0.659 0.631 0.629 1.328 0.865 1.34 2.19 720 0.721 0.656 0.731 0.669 0.729 0.661 0.878 0.738 0.723 0.650 1.489 0.927 1.10 0.76 Electricity 96 0.131 0.228 0.149 0.252 0.136 0.232 0.190 0.301 0.189 0.298 1.528 0.927 3.68 1.72 192 0.141 0.241 0.160 0.264 0.167 0.269 0.202 0.313 0.204 0.311 1.521 0.927 11.87 8.71 336 0.153 0.260 0.170 0.279 0.169 0.275 0.210 0.318 0.231 0.336 1.530 0.934 9.47 5.45 720 0.171 0.282 0.204 0.314 0.200 0.298 0.239 0.350 0.246 0.361 1.535 0.942 14.5 5.37 Weather 96 0.178 0.261 0.169 0.251 0.175 0.259 0.364 0.406 0.319 0.329 0.258 0.254 -5.33 -3.98 192 0.234 0.315 0.229 0.300 0.225 0.299 0.436 0.439 0.322 0.354 0.339 0.294 -4.00 -7.14 336 0.381 0.443 0.321 0.403 0.352 0.412 0.470 0.469 0.401 0.391 0.421 0.343 -18.69 -29.15 720 0.549 0.536 0.414 0.465 0.421 0.481 0.467 0.447 0.433 0.444 0.465 0.394 -32.60 -36.04
-Table 2: Experimental results on real data in multivariate input, multivariate output setting.
-The results were averaged over five runs to eliminate the effect of randomness.
-The IMP column displays the error reduction of our proposed model compared to the best baseline.
-The best results are highlighted in bold and the second-best results are in underlined.
-Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.864 0.703 0.874 0.699 0.872 0.701 0.947 0.714 0.914 0.695 1.579 0.869 0.92 -0.57 192 0.903 0.719 0.926 0.728 0.931 0.732 1.008 0.744 0.973 0.739 1.725 0.926 2.48 1.24 336 0.916 0.722 0.954 0.738 0.955 0.745 1.057 0.792 0.969 0.744 1.799 0.948 3.98 2.17 720 0.945 0.735 0.986 0.746 1.015 0.783 1.063 0.814 1.009 0.751 1.877 0.977 4.16 1.47 Carbon 96 0.232 0.370 0.235 0.375 0.244 0.381 0.731 0.858 0.418 0.561 1.771 1.039 1.28 1.33 192 0.248 0.389 0.273 0.406 0.284 0.412 0.779 0.871 0.445 0.594 1.853 1.077 9.16 4.19 336 0.266 0.410 0.307 0.430 0.305 0.427 0.801 0.896 0.457 0.612 1.919 1.106 12.79 3.98 720 0.282 0.425 0.368 0.476 0.364 0.469 0.818 0.914 0.551 0.687 2.025 1.141 22.53 9.38 Electricity 96 0.148 0.273 0.166 0.292 0.169 0.293 0.194 0.309 0.200 0.291 1.221 0.858 10.84 6.19 192 0.176 0.298 0.189 0.312 0.191 0.320 0.211 0.319 0.209 0.319 1.223 0.857 6.87 4.49 336 0.188 0.310 0.209 0.329 0.204 0.321 0.218 0.330 0.211 0.334 1.297 0.881 7.84 3.43 720 0.214 0.335 0.252 0.371 0.255 0.374 0.240 0.351 0.258 0.359 1.369 0.908 10.83 4.56 Weather 96 0.096 0.206 0.127 0.234 0.131 0.237 0.109 0.212 0.112 0.243 0.130 0.262 11.93 2.83 192 0.378 0.416 0.412 0.425 0.423 0.427 0.401 0.419 0.439 0.461 0.451 0.478 5.74 0.72 336 0.780 0.604 0.834 0.613 0.845 0.621 0.806 0.605 0.812 0.604 0.844 0.620 3.23 0.00 720 1.237 0.797 1.408 0.825 1.392 0.818 1.398 0.816 1.418 0.876 1.470 0.934 11.14 2.33
-Table 3: Experimental results on real data in multivariate input, univariate output setting.
-The results were averaged over five runs to eliminate the effect of randomness.
-The IMP column displays the error reduction of our proposed model compared to the best baseline.
-The best results are highlighted in bold and the second-best results are in underlined.
-Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.867 0.695 0.874 0.699 0.872 0.701 0.961 0.732 0.917 0.699 1.579 0.869 0.57 0.57 192 0.921 0.724 0.926 0.728 0.931 0.732 1.023 0.760 0.971 0.741 1.725 0.926 0.54 0.55 336 0.950 0.736 0.954 0.738 0.955 0.745 1.084 0.803 1.001 0.747 1.799 0.948 0.42 0.27 720 0.984 0.747 0.986 0.746 1.015 0.783 1.092 0.833 1.013 0.755 1.877 0.977 0.20 -0.13 Carbon 96 0.216 0.357 0.235 0.375 0.244 0.381 0.615 0.796 0.422 0.567 1.771 1.039 8.08 4.80 192 0.264 0.398 0.273 0.406 0.284 0.412 0.631 0.818 0.446 0.601 1.853 1.077 3.29 1.97 336 0.293 0.422 0.307 0.430 0.305 0.427 0.722 0.846 0.463 0.615 1.919 1.106 3.3 1.17 720 0.360 0.471 0.368 0.476 0.364 0.469 0.779 0.875 0.557 0.690 2.025 1.141 1.10 -0.42 Electricity 96 0.154 0.277 0.166 0.292 0.169 0.293 0.198 0.311 0.209 0.291 1.221 0.858 7.23 4.81 192 0.173 0.297 0.189 0.312 0.191 0.320 0.214 0.323 0.223 0.342 1.223 0.857 8.46 4.81 336 0.196 0.315 0.209 0.329 0.204 0.321 0.231 0.348 0.231 0.345 1.297 0.881 3.92 1.87 720 0.238 0.357 0.252 0.371 0.255 0.374 0.288 0.389 0.307 0.398 1.369 0.908 5.55 3.77 Weather 96 0.119 0.228 0.127 0.234 0.131 0.237 0.138 0.245 0.157 0.282 0.130 0.262 6.30 2.56 192 0.421 0.433 0.412 0.425 0.423 0.427 0.476 0.501 0.470 0.483 0.451 0.478 -2.18 -1.88 336 0.843 0.621 0.834 0.613 0.845 0.621 0.864 0.658 0.872 0.684 0.844 0.620 -1.07 -1.30 720 1.414 0.826 1.408 0.825 1.392 0.818 1.412 0.835 1.555 0.954 1.470 0.934 -1.58 -0.97
-Table 4: Experimental results on real data in univariate input, univariate output.
-The results were averaged over five runs to eliminate the effect of randomness.
-The IMP column displays the error reduction of our proposed model compared to the best baseline.
-The best results are highlighted in bold and the second-best results are in underlined.
-As shown in Tables 2, 3, and 4, our proposed model has achieved the best or near-best results across all settings.
-We present detailed results in three different settings to ensure a fair comparison of the strengths and weaknesses of the baseline models and to clearly explain the effectiveness of our proposed model.
+We consider 4 long-term forecast horizons, namely 96, 192, 336, and 720, as they are commonly used in long-term time-series forecasting studies Wu et al.
+(2021); Zhou et al.
+(2022); Zeng et al.
+(2023).
+These multi-day to multi-week horizons are also more suitable for downstream applications, such as generator scheduling, carbon-aware computing Maji et al.
+(2022), and city-scale planning Mouatadid et al.
+(2024), where long-range forecasts support critical decision-making tasks more effectively than short-term ones.
+Dataset Look-ahead (# samples) DST s-Mamba DLinear Autoformer TimesFM PatchTST Repeat Last MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.864 0.703 0.874 0.699 0.872 0.701 0.947 0.714 0.914 0.695 0.901 0.699 1.579 0.869 192 0.903 0.719 0.926 0.728 0.931 0.732 1.008 0.744 0.973 0.739 0.944 0.735 1.725 0.926 336 0.916 0.722 0.954 0.738 0.955 0.745 1.057 0.792 0.969 0.744 0.958 0.753 1.799 0.948 720 0.945 0.735 0.986 0.746 1.015 0.783 1.063 0.814 1.009 0.751 0.989 0.803 1.877 0.977 Carbon 96 0.232 0.370 0.235 0.375 0.244 0.381 0.386 0.506 0.418 0.561 0.349 0.497 1.771 1.039 192 0.248 0.389 0.273 0.406 0.284 0.412 0.371 0.491 0.445 0.594 0.413 0.569 1.853 1.077 336 0.266 0.410 0.307 0.430 0.305 0.427 0.454 0.532 0.457 0.612 0.435 0.588 1.919 1.106 720 0.282 0.425 0.368 0.476 0.364 0.469 0.407 0.502 0.551 0.687 0.512 0.617 2.025 1.141 Electricity 96 0.148 0.273 0.166 0.292 0.169 0.293 0.194 0.309 0.200 0.291 0.198 0.311 1.221 0.858 192 0.176 0.298 0.189 0.312 0.191 0.320 0.211 0.319 0.209 0.319 0.214 0.326 1.223 0.857 336 0.188 0.310 0.209 0.329 0.204 0.321 0.218 0.330 0.211 0.334 0.237 0.349 1.297 0.881 720 0.214 0.335 0.252 0.371 0.255 0.374 0.240 0.351 0.258 0.359 0.284 0.397 1.369 0.908 Weather 96 0.096 0.206 0.127 0.234 0.131 0.237 0.109 0.212 0.112 0.243 0.115 0.256 0.130 0.262 192 0.378 0.416 0.412 0.425 0.423 0.427 0.401 0.419 0.439 0.461 0.437 0.459 0.451 0.478 336 0.780 0.604 0.834 0.613 0.845 0.621 0.806 0.605 0.812 0.605 0.829 0.618 0.844 0.620 720 1.237 0.797 1.408 0.825 1.392 0.818 1.398 0.816 1.418 0.876 1.378 0.133 1.470 0.934
+Table 2: Comparison of the forecasting accuracy of DST and baselines on four dataset in the multivariate input, univariate output setting.
+The results were averaged over five runs to mitigate the effect of randomness.
+The best results are highlighted in bold and the second-best results are underlined.
+As shown in Table 2, DST achieves the best performance, with respect to MSE and MAE, in almost all settings.
+Notably, it outperforms the best baseline model, which is s-Mamba in most cases, by up to 9.10% on average.
+It can be seen that the gap between DST and the second best method generally widens for longer forecast horizons.
+This highlights the strength of the proposed framework for long-term multivariate time-series forecasting.
 Figure 6: MSE comparison of our proposed model against baselines on different datasets.
@@
 Note that the axis scales are exaggerated for better visualization of the differences between models.
-DLinear and NLinear achieve the best performance among the baselines.
-Despite their impressive performance, they are univariate models that lack a mechanism to extract spatial dependencies.
+To further demonstrate the effectiveness of DST, we provide a visual comparison of its predictions against those of the best-performing baselines, s-Mamba and DLinear, in Figure 7.
+These examples highlight the ability of our model to more accurately capture the temporal dynamics of the target time-series across different domains.
+In particular, DST demonstrates smoother forecasts and better alignment with the ground truth values.
+These qualitative results complement the quantitative improvements reported in Table 2, offering additional insight into the predictive strengths of our approach.
+(a) Air Pollution
+(b) Carbon Intensity
+(c) Electricity
+(d) Weather
+Figure 7: Comparison between the forecasts of DST, and two best performing baselines: s-Mamba and DLinear, on selected samples from a randomly chosen variate in each dataset.
+Apart from s-Mamba, DLinear also exhibits reasonable performance among the baselines.
+Despite its good performance, it is a univariate model that lacks a mechanism to extract spatial dependencies.
 This limitation becomes evident in the multivariate input and univariate forecasting setting, where forecasting a variate may benefit from capturing lagged dependencies with other variates.
-Univariate models like DLinear or NLinear cannot leverage this information, leading to inferior performance compared to models that can, such as our proposed model and transformer-based models.
-This is particularly evident in the forecasting of PM2.5 with multivariate input.
+Univariate models cannot leverage this information, leading to inferior performance compared to the models that can, such as our proposed model (DST) and transformer-based models.
+This is particularly evident in the forecasting of PM2.5 air pollution dataset.
 Transformer-based forecasting models, such as Autoformer, Fedformer, and Informer, are highly complex and over-parameterized, making them susceptible to overfitting.
-This issue is apparent in the forecasting of the carbon dataset.
-Additionally, we have included the inference results of TimesFM, which were obtained without additional training or fine-tuning.
-This lack of fine-tuning likely contributes to its suboptimal performance, as it might not have adapted well to the specific characteristics of the datasets.
-Pretrained models are generally not tailored to the nuances of new datasets unless they undergo further training to update their parameters.
-Consequently, the model’s inability to learn and adjust to the unique temporal patterns and data distribution of the datasets has likely resulted in less accurate forecasts compared to the other models that were specifically trained.
-Figure 8 presents a comparative analysis of the Mean Squared Error (MSE) performance of our proposed model, DST, against several baseline models across four datasets: Beijing PM2.5, Carbon, Electricity, and Weather.
+This issue is apparent in the forecasting of carbon intensity.
+Additionally, we have included the forecasting performance of TimesFM, which was obtained without fine-tuning (i.e.
+its zero-shot performance, which was also reported in Das et al.
+(2024)).
+This lack of fine-tuning likely contributes to its subpar performance, as it might not have adapted well to the specific characteristics of each dataset.
+Pretrained models must typically undergo additional training to align with the unique temporal patterns and distributions of new datasets.
+Without this adaptation, performance may suffer in comparison to models trained directly on the target datasets.
+Figure 6 presents a comparative analysis of the forecasting performance, in terms of the distribution of MSE, of our proposed model, DST, against several baseline models across four datasets: Air pollution, Carbon, Electricity, and Weather.
+The observed variability in MSE across runs primarily stems from stochastic elements in the training process, including random weight initialization, and data shuffling due to mini-batch sampling.
+One baseline model, TimesFM, is excluded from the plots because it is evaluated solely in the zero-shot inference setting.
+Consequently, it produces the MSE value that is reported in Table 2, with no variability.
 This evaluation is conducted on a multivariate input-multivariate output forecasting task, where the input and output lengths are set to 336 and 96, respectively.
 The results demonstrate that DST consistently achieves lower MSE values compared to the baselines, indicating its superior forecasting capability.
-Notably, traditional transformer-based models, such as Autoformer and Informer, tend to exhibit higher variance and median errors, suggesting increased instability in their predictions.
-DLinear, a linear baseline, performs competitively but is outperformed by DST, particularly on the Carbon dataset.
-The y-axis scales are adjusted for better visualization of performance differences among the models
-6 Ablations
-To thoroughly examine the properties of our proposed model, we conducted ablation studies to evaluate the effectiveness of each of its key modules.
-The modules under evaluation include (1) Decomposition, (2) Date and Time Embedding, (3) Temporal Feature Extraction, and (4) Spatial Feature Extraction.
-For each dataset, we compare the model’s performance with and without each of these modules.
-Due to the modular and pipelined structure of the proposed architecture, we can easily remove individual modules without impacting the functionality of the remaining components.
+Transformer-based models, such as Autoformer and PatchTST, exhibit relatively higher median errors and greater variability.
+DLinear and s-Mamba perform relatively well, but are still outperformed by DST in all datasets.
+We note that the y-axis is exaggerated in some cases to better highlight performance differences between models.
+6 Ablation Study
+We conducted an ablation study to evaluate the effectiveness of each module in DST to better explain its performance.
+The modules under evaluation are (1) Decomposition, (2) Date and Time Embedding, (3) Temporal Feature Extraction, and (4) Spatial Feature Extraction.
+For each dataset, we compare the model’s performance without each of these modules.
+Due to the modular structure of DST, we can easily remove individual modules without impacting the functionality of the remaining components.
 The only exception is the Date and Time Embedding module, whose output is concatenated with the extracted spatio-temporal features.
-To simulate the removal of this module, we replace its output with zero-vectors, thereby mitigating its effect on the model’s performance.
+To remove the influence of this module, we replace its output with a zero vector.
 Table 3 demonstrates the effectiveness of each module within the model architecture.
-Among all the components, the temporal feature extraction module plays the most influential role, as indicated by the highest increase in Mean Squared Error (MSE) when this module is removed.
+Among all the components, the temporal feature extraction module plays the most influential role, as indicated by the highest increase in MSE when this module is removed.
 This suggests that capturing temporal dependencies is crucial for accurately modeling the datasets, particularly for data characterized by strong temporal patterns, such as air pollution, carbon emissions, and electricity usage.
-The ablation results also reveal that other modules, such as decomposition, date and time embedding, and spatial feature engineering, contribute significantly to model performance, albeit to a lesser extent.
+Decomposition is also among the most effective components, especially for air pollution and weather forecasting.
+The ablation results also reveal that other modules, i.e.
+date and time embedding and spatial feature extraction, contribute significantly to model performance, albeit to a lesser extent.
 These components work synergistically to enhance predictive accuracy by providing complementary information that captures different aspects of the data, such as spatial correlations and temporal variations.
 Overall, the study reveals the necessity of taking a multi-faceted approach that integrates various techniques to effectively handle complex, real-world datasets.
-Additionally, we conducted experiments to evaluate the impact of the number of GATv2 layers.
-Across all experiments, a single layer for spatial feature extraction consistently yielded the best performance.
-Figure 9 illustrates how the MSE changes as more GATv2 layers are added for the carbon intensity dataset.
+Additionally, we carried out experiments to evaluate the impact of the number of GATv2 layers on DST’s performance.
+We found that across all experiments, a single layer for spatial feature extraction consistently yields the best performance.
+As a result, DST is implemented with a single GATv2 layer.
 Dataset Air Pollution Carbon Electricity Weather
-w/o Decomposition 0.441 0.474 0.154 0.179
-w/o Date&Time Emb. 0.424 0.482 0.149 0.158
-w/o Temporal F.E. 0.452 0.491 0.162 0.161
-w/o Spatial F.E. 0.439 0.488 0.156 0.165
-Original 0.404 0.460 0.133 0.155
+DST w/o Decomposition 0.441 0.474 0.154 0.179
+DST w/o Date&Time Emb. 0.424 0.482 0.149 0.158
+DST w/o Temporal F.E. 0.452 0.491 0.162 0.161
+DST w/o Spatial F.E. 0.439 0.488 0.156 0.165
+DST 0.404 0.460 0.133 0.155
 Table 3: Ablation studies on the effectiveness of key modules. Reported values are the MSE of each setting.
 7 Conclusion
-In this paper, we introduced a novel multivariate time-series forecasting model that leverages the advanced capabilities of GATv2 to capture spatial dependencies among different time-series variables.
-Our approach, which incorporates a decomposition-based preprocessing step, has demonstrated incremental improvements in forecast accuracy and reliability across various real-world datasets.
-The experimental results underscore the potential of our model to outperform existing state-of-the-art methods, particularly in complex domains such as energy management, weather prediction, carbon intensity forecasting, and air pollution monitoring.
-By effectively modeling the intricate interdependencies within multivariate time-series data, our proposed model provides a robust framework for long-term forecasting, contributing to the development of smart infrastructure systems and enhancing societal outcomes.
-Future work will focus on further optimizing the model, exploring its applicability to additional domains, and integrating adaptive graph learning techniques to enhance its performance in dynamic environments.
-Appendix
-Appendix A Number of GATv2 layers
-Figure 9: Changes of MSE loss by increasing the number of GATv2 layers used for spatial feature extraction on carbon intensity dataset.
+In this paper, we introduced a novel multivariate time-series forecasting model that leverages a GNN architecture integrating GATv2 and TCN.
+This GNN relies on a distinct graph structure per time-series component.
+To learn this graph structure, we implemented a preprocessing step to decompose the time-series into its fundamental components and proposed an effective algorithm for learning the per-component graph structure, based on historical sensor data.
+Extensive experiments on four real-world urban datasets with different temporal resolutions, numbers of sensors, and interdependencies confirm that DST consistently achieves the best long-term forecasting performance across different forecast horizons.
+Specifically, it outperforms the best-in-class time-series forecasting model by 2.89% to 9.10% in different datasets.
+Future work will focus on using these forecasting models in various decision-making tasks in the urban environment, from infrastructure planning to assignment of AI workloads (particularly training and fine-tuning) to datacenters, to quantify improvements in decision quality resulting from more accurate long-term forecasts.
+Furthermore, we intend to explore how training the long-term forecasting models using the task loss would change the decision quality.